### v1

# Word length predicts word order: “Min-max”-ing drives language evolution

Abstract
Current theories of language propose an innate (Baker 2001; Chomsky 1981) or a functional (Greenberg 1963; Dryer 2007; Hawkins 2014) origin for the surface structures (i.e. word order) that we observe in languages of the world, while evolutionary modeling (Dunn et al. 2011) suggests that descent is the primary factor influencing such patterns. Although there are hypotheses for word order change from both innate and usage-based perspectives for specific languages and families, there are key disagreements between the two major proposals for mechanisms that drive the evolution of language more broadly (Wasow 2002; Levy 2008). This paper proposes a universal underlying mechanism for word order change based on a large tagged parallel dataset of over 1,500 languages representing 133 language families and 111 isolates. Results indicate that word class length is significantly correlated with word order crosslinguistically, but not in a straightforward manner, partially supporting opposing theories of processing, while at the same time predicting historical word order change in two different phylogenetic lines and explaining more variance than descent or language area in regression models. Such findings suggest an integrated “Min-Max” theory of language evolution driven by competing pressures of processing and information structure, aligning with recent efficiency-oriented (Levshina 2023) and information-theoretic proposals (Zaslavsky 2020; Tucker et al. 2025).
1 Introduction
Theories of language should describe not only the current state of languages, but also explain how they have developed (their history) and predict how they will change (their future). To date, however, existing theories have primarily focused on describing the present state of language, proposing that the current diversity in surface structures or patterns that we observe in languages of the world, such as for word order, have either an innate (Baker 2001; Chomsky 1981) or functional (Greenberg 1963; Dryer 2007; Hawkins 2014) source. More recent diachronic work has tried to answer the second question on how languages develop. From an “innateness” perspective, surface word orders are seen to be derived from an underlying (universally shared) structure, whereby parameters for surface realizations change over time (Carnie & Guilfoyle 2000; Crisma & Longobardi 2009; Martins & Cardoso 2018), while from a “functional” perspective surface word orders are the result of cumulative changes in linked structures (sometimes called ‘constructions’) that compete for adoption in individual languages (see Gerritsen & Stein 1992; Leino & Kuningas 2006; Barðdal & Gildea 2016).
Importantly, none of these perspectives are particularly explanatory of drivers in language evolution that affect all languages similarly. The closest observations in this regard come from evolutionary modeling, where descent from a common ancestor was found to be the most explanatory factor in predicting word order (Dunn et al. 2011; though see Jäger & Wahle 2021; Naranjo & Becker 2022 for alternate perspectives), and language processing, where word length (Piantadosi et al. 2011; Wasow 2022) and surprisal (Levy 2008) have been shown to impact the realization of word order. There are also more general claims that “efficiency” (variously defined) explains grammatical (Hawkins 2014; Levshina 2023) or semantic (Zaslavsky 2020) change. However, the fact that languages tend to display word orders like their family members does not explain the variation in word orders that we find within families, and perspectives on processing make very different predictions: the literature on word length predicts that short/light words will occur first in a sentence, while the literature on surprisal predicts that long/heavy words will occur first. Further, the literature on efficiency does not make actual testable cross-linguistic predictions about basic word order.
There are also other issues to consider, based on observation of the languages for which we have data. Typological surveys (Dryer & Haspelmath 2013; Skirgård et al. 2023; Bickel et al. 2023) reveal that over 80% of the sampled languages have a Subject-initial basic word order, while only about 10% have a Verb-initial order. Further, there are various correlations to be found between the orders of clausal elements, such that basic word order is related to how other clause types are ordered. While all languages allow variation in word order to different degrees, the word order most often found in declarative sentences can predict (to some degree) the ordering of other grammatical elements in a language (Dryer 1992, 2011).
It is not completely clear why one basic constituent order correlates better with one particular order of (e.g.) noun phrase elements than another, nor why some basic word orders are less well-represented crosslinguistically. This raises the question whether there are underlying processes that drive these observed tendencies. Some explanation clearly comes from descent, i.e. that languages tend to resemble their ancestral forms. However, the word order changes that languages undergo over time suggests other pressures in terms of how complex linguistic elements are and how these constituents are processed in the brain (Maurits & Griffiths 2014; Schouwstra & de Swart 2014, see also Piantadosi et al. 2011; Bickel et al. 2015; Ding et al. 2017; Fernandez 2016; Jaeger & Tily 2011; Tily 2010; Sinnemäki 2014).
Answers to these questions have been elusive in part due to a lack of data that would facilitate crosslinguistic investigations. This is because such investigations require data that is annotated for various word types or dependencies, which is unavailable for the majority of the world’s languages. The largest sample of languages used to investigate these questions looked at annotated corpora in 80 languages (Hahn & Xu 2022), the majority of which are from the Indo-European language family. Such data risks biasing findings toward a particular group of languages (Blasi et al. 2022), in much the same way that other scientific findings have historically been biased toward WEIRD demographics (Henrich et al. 2010), leading to recent calls for more expansive datasets to examine linguistic questions (Benítez-Burraco 2025; Levshina et al. 2023).
The current study addresses this by developing a large corpus of texts tagged for word class. With data sourced from the Parallel Bible Corpus (PBC; Mayer & Cysouw 2014), machine learning techniques allow for automatic tagging [^1] Automated part of speech tagging was validated via comparison with existing taggers for high-resource languages (SpaCy, 16 languages - avg nouns: 0.83, verbs: 0.74; Trankit, 31 languages - avg nouns: 0.83, verbs: 0.76) and gold-standard tagged data in the Universal Dependencies Treebanks (UDT; Zeman et al. 2024; 106 languages; https://universaldependencies.org). See Ring (2025) for more details. of word classes in 1,597 languages (roughly 20% of the world’s languages) that represent 133 families and 111 isolates.222‘Isolates’ being defined here as languages in the corpus that share no known family membership with other languages in the corpus. Individual corpora in the majority of these languages (over 1,500) contain more than 1,800 parallel sentences, and the smallest corpus contains more than 700 sentences. [^3] For clarity, the UDT has 160+ hand-annotated languages available for crosslinguistic investigation, but these are not parallel texts and the number of sentences in each language is variable, with some having fewer than 200 sentences in the v2.14 release. The ODIN project (and its re-packaged form via LECS: https://huggingface.co/datasets/lecslab/glosslm-corpus) has interlinear glossed text (IGT) from over 1,800 languages, but the majority of these are single-word/line examples, and only 300 of the languages in their dataset have more than 100 sentences. Additional cross-linguistic investigations using tagged corpora include Jing et al. (2021: 71 languages) and Hahn et al. (2020: 51 languages). This dwarfs both the linguistic data and the language diversity available from other sources and means that crosslinguistic investigations using this dataset can be confident that they are comparing similar contexts. The dataset (the tagged PBC) and explanations regarding its development, along with Python code for replicating the results in this paper, can be found at the linked repository [^4] https://github.com/lingdoc/taggedPBC and is described at greater length in Ring (2025).
This dataset can be used to investigate the crosslinguistic questions highlighted above regarding whether there is some pressure (beyond descent) that explains the evolution of word order in languages of the world. If there is indeed a mechanism that underlies the observed surface structures in present-day languages, this should also be one that reveals historical processes and allows us to predict future developments. Specifically, by deriving a novel corpus-based measure of word order (the “N1 ratio”) [^5] The N1 ratio was validated via comparison with existing classifications of languages in the tagged PBC and three typological databases: the World Atlas of Language Structures (WALS; Dryer & Haspelmath 2013; https://wals.info/), the AUTOTYP Database (Bickel et al. 2023; https://github.com/autotyp/autotyp-data), and Grambank (Skirgård et al. 2023; https://grambank.clld.org/) where the N1 ratio significantly differentiated (p ¡ 0.001) between SV/VS/free classifications in all three databases. See Ring (2025) for more details. from this dataset, the current paper investigates the relationship between basic word order and the lengths of particular word classes (nouns/arguments and verbs/predicates) crosslinguistically. Results show that the lengths of these word classes is significantly related to word order, that they can predict historical word order change in two different lineages, and that they account for more variance than family membership (descent) or areal patterns (geography) when examined using hierarchical linear regressions. Such findings suggest the possibility for a unified (“Min-Max”) theory of language evolution that can explain both observed surface structures in languages of the world as well as how such patterns have developed diachronically.
2 Complexity, processing, and word order
Linguistic complexity can take many forms, with many considerations regarding its measurement (see Jaeger & Tily 2011). As noted by Joseph and Newmayer (2012: 341):
“how linguistic complexity is conceived of, and how it might be identified and measured, has itself never been agreed upon by linguists. It cannot be neatly separated from theoretical considerations of where language structure ends and language processing (or use) begins, or where variation (whether social or stylistic) is located.”
Some current efforts to measure linguistic complexity focus on quantifying the overall complexity of a language (Bentz et al., 2016, 2017; Baerman et al., 2015, 2017), and such measures may be under considerable debate (see Bentz et al., 2023). Additional research suggests that word complexity can be measured by word length (see Lewis & Frank, 2016; and see Piantadosi et al., 2011; Meylan & Griffiths, 2021; Koplenig et al., 2022 for further discussion in relation to frequency in corpora). Words which are formed from longer sequences of phonemes are more difficult to process, and this is also related to their conceptual complexity. The fact that word length is related to a word’s complexity is important for our understanding of word order, i.e. how closely is the ordering of constituents related to how we process words/structures of varying lengths?
Research into how speakers process words has identified several constraints and tendencies that have various implications for how languages develop. Hawkins (2014) summarizes this literature in arguing for a comprehensive theory (the Performance-Grammar Correspondence Hypothesis; PGCH) based firmly in the idea of efficiency as an explanation for a large number of observed patterns and historical changes cross-linguistically. [^6] Levshina (2023) argues for another explanation termed “communicative” efficiency, which identifies “accessibility” as a factor in language use. This seems to overlap to some degree with the effects identified by “surprisal”. As Levshina’s observations do not make specific predictions about word order crosslinguistically, I revisit them in the discussion. His essential claim is that speakers minimize complexity (in forms, distance) in order to speed up processing, a claim that aligns with observations about constituent “weight” (Wasow, 2002; Zipf, 1949) and “dependency length minimization” (Futrell et al., 2015). However, this claim contrasts with findings from Surprisal Theory (Hale 2001; Levy 2008) that speakers maximize complexity, i.e. that more complex forms are preferred by speakers since they support information content/retrieval and reduce entropy in utterances (see Rezaii et al. 2023).
These two opposing perspectives on processing (the minimization vs maximization of complexity) make two different predictions for word order crosslinguistically. The PGCH claim (and Wasow’s) is that “heavier” or more complex constituents are typically “put off” until later in the sentence, while less complex constituents are put first. This predicts that a language in which nouns in general are easier to process than verbs should display SV order, while one in which verbs are easier to process should display VS order. Yet evidence from languages like Japanese, which places heavier phrases closer to the beginning of the sentence (Hawkins, 1994; Yamashita & Chang, 2001; Yamashita, 2002; and see discussion in Wasow, 2022) supports the opposing claim by Surprisal Theory that words with higher information content (or which improve predictability of following words) are more likely to occur first. Here the prediction is that a language with heavier nouns should display SV order, and one with heavier verbs should display VS order.
Such predictions, if true, should be observable in crosslinguistic corpora. An important note here is that while the processing literature is largely grounded in a functional/usage-based perspective, it does not necessarily require a usage-based stance. That is, cognitive processes can be argued to apply at various levels of processing, including above the “innate” bedrock of current Minimalist theory (see Rastelli 2025). If there is a universal processing bias in the brain related to the complexity of structures or words in language, as has been claimed, such a bias would apply regardless of whether grammatical rules are innate or learned, and would impact or drive how such rules are realized or change over time.
2.1 Main hypotheses
Using word length as a measure of word complexity, the current paper compares average lengths of particular classes of words (nouns/arguments and verbs/predicates) with word order in individual languages. The two processing claims in the literature give rise to the following two competing hypotheses regarding this relationship:
H1:
SV languages have shorter “S” than “V”, while VS languages have shorter “V” than “S”.
The average length of nouns is shorter in languages classed as Subject-initial, while the average length of verbs is shorter in languages classed as Verb-initial.
(minimize complexity)
H2:
SV languages have longer “S” than “V”, while VS languages have longer “V” than “S”.
The average length of nouns is longer in languages classed as Subject-initial, while the average length of verbs is longer in languages classed as Verb-initial.
(maximize complexity)
2.2 Investigating word order and length
Using the 1,597 languages in the tagged PBC, we can identify basic word order via the N1 ratio. [^7] As an example, English [ISO 639-3: eng] has an N1 ratio of 4.51 and is classed as SV, whereas Irish [ISO 639-3: gle] has an N1 ratio of 1.24 and is classed as VS. Details on the N1 ratio and its validity are given above and at the linked repository (Fn 4). We can also identify the average length of nouns/arguments and verbs/predicates in each language’s corpus and see how well this correlates with the N1 ratio.
2.2.1 Word order
Of the languages in the tagged PBC, 976 are also found in typological databases with categorizations for word order. For the remainder, the work described at the linked repository has shown that the N1 ratio is a valid corpus-based proxy for word order. Accordingly I used the N1 ratio to impute the (SV, VS, free) category of the remaining 621 languages in the tagged PBC not categorized by WALS, AUTOTYP or Grambank. [^8] I trained a Gaussian Naive Bayes classifier on the known data to predict the category (SV, VS, free) of languages for which word order status was unknown, based on the N1 ratio as derived from the PBC. Code to reproduce these results can be found at the linked repository (Fn 4). This resulted in 1,251 languages classified as “SV”, 295 languages classified as “VS”, and 51 languages classified as “free”.
2.2.2 Nouns and verbs
To measure word length, I counted the number of characters in a word tagged as Noun/Argument or Verb/Predicate and computed an average length for each class on a per-language basis. [^9] For example, English [ISO 639-3: eng] arguments have an average length of 5.9 characters, while English predicates have an average length of 5.75. Irish [ISO 639-3: gle] arguments have an average length of 6.31 characters, while Irish predicates have an average length of 6.29 characters. In the tagged PBC corpus the average number of (unique) arguments per language is 732, while the average number of predicates is 495. The mean length of arguments is 6.76 characters, while the mean length of predicates is 6.90 characters. Conducting a Student’s T-Test showed that this is statistically significant (p = 0.006), indicating that crosslinguistically, arguments are shorter than predicates. All statistics showed a bimodal (normal) distribution.
2.2.3 Relationship between word order and noun/verb lengths
To see whether word order is related to the length of nouns and verbs I conducted a repeated-measures ANOVA with the average lengths of nouns and verbs as the repeated measures and the word order (SV, VS, free) as the between subject factors. Analysis of the results of word order and its relation to length of the different argument and predicate categories is quite striking. First, both SV and VS languages have a highly significant effect of word length, [^10] Additional statistical comparisons and plots are available at the linked repository (Fn 4). whereby arguments are, overall, shorter than predicates crosslinguistically (Figure 1; p < 0.001). [^11] It is also worth noting here that in this corpus words are overall shorter in SV languages than VS languages, which in turn are shorter than those in ‘free’ languages. This displays a general observed trend that languages with free word order tend to have more morphology to identify argument roles and other grammatical relations (Koplenig et al. 2017; Yadav et al. 2020).
Figure 1: Normalized lens of Args, Preds, by word order
Figure 2: Frequency-weighted lens of Ns, Vs, by word order
If we consider only nouns and only verbs, the findings are similar when we normalize the dataset by reducing the vocabulary to unique items. However, if we allow for frequency effects (i.e. we do not reduce the vocabulary to unique items) there is a clear correlation with word order: nouns are longer than verbs in SV languages, but shorter than verbs in VS languages (Figure 2; p < 0.001). This is a major finding that we will return to in the discussion. Setting it aside for now, we turn to two different tests of this measure relating to diachrony: predicting word order change, and the potential interaction with family membership (descent) and area (geography).
2.3 Testing the predictive validity of word length
Given the observation that relative lengths of nouns and verbs differ in corpora for languages with different word orders, we can ask how predictive this observation is. That is, can we predict the word order of a language based on the relative lengths of its nouns and verbs? If we are strictly interested in identifying basic word order from corpora, it is clear that the N1 ratio already allows us to do so. However, we are also interested in understanding how noun and verb lengths may be related to word order. Being able to predict word order from noun and verb lengths in a corpus for both a modern language and its ancestral language would indicate that such properties are related to how word order evolves across time.
To investigate this question, I sourced additional corpora of tagged data from two pairs of historical and modern languages in which the modern language changed its word order: 1) Ancient Hebrew (VS) and Modern Hebrew (SV) and 2) Classical Arabic (VS) and Egyptian Arabic (SV). All languages belong to the Semitic branch of the Afroasiatic language family, a necessary constraint given the limited amount of data available for most pre-modern languages. In the case of Hebrew, the UDT has POS-tagged sentences for both Ancient [^12] https://universaldependencies.org/treebanks/hbo_ptnk and Modern [^13] https://universaldependencies.org/treebanks/he_htb varieties. In the case of Classical Arabic I used the Quranic Arabic Corpus [^14] Version 0.4: https://corpus.quran.com and for Egyptian Arabic I used the BOLT Egyptian Arabic Treebank (Maamouri et al. 2018).151515(Discussion Forum) https://catalog.ldc.upenn.edu/LDC2018T23 If the lengths of nouns and length of verbs in these corpora allows us to correctly classify Ancient Hebrew and Classical Arabic as VS and the modern varieties as SV, this would suggest that these are salient properties affecting how word order has developed diachronically. Importantly, given the textual differences between the tagged PBC (New Testament texts) and the test corpora (Biblical, Quranic, modern, and informal texts), such a finding would be significant.
In order to conduct this test, the data for Ancient Hebrew (ISO 639-3: hbo), Modern Hebrew (ISO 639-3: heb), Classical Arabic (cla) [^16] Classical Arabic does not have an assigned ISO 639-3 code or Glottocode. Accordingly, I use the abbreviation cla here, as there is no current language assigned to this code in the ISO 639-3 tables. and Egyptian Arabic (ISO 639-3: arz) were processed and the respective word classes were counted and measured as per the method outlined for the tagged PBC (each corpus > 400 sentences, see Table 1). The average length of nouns and verbs (including frequency effects) were then used as continuous features for a Gaussian Naive Bayes (GNB) classifier.
Language ISO 639-3 Data Source Sentences Unique nouns Unique verbs
Ancient Hebrew hbo UDT v2.14 409 734 782
Modern Hebrew heb UDT v2.14 882 2068 1297
Classical Arabic cla Quran 6,236 8,243 7,966
Egyptian Arabic arz BOLT 31,688 16,698 10,581
Table 1: Corpora used for testing historical word order change
Three additional features were also used to expand the feature dimensions. The first is a categorical feature: whether a language’s nouns are longer than its verbs. The second is a continuous feature: the difference in length between nouns and verbs. This second measure is related to minimization of complexity, based on the idea that a larger difference between lengths of word classes may effect the likelihood of that particular class being placed first. The third measure is the ratio of noun and verb lengths, and is related to maximization of complexity (via “entropy”), whereby a lower ratio indicates that the particular word classes are less likely to be easily differentiable (and therefore predictable) based on their length.
Using these five features, a GNB classifier was trained on the original combined word order dataset containing hand-coded classification of 974 languages (Modern Hebrew and Modern Arabic were removed), and then used to predict word order for our pairs of languages: Ancient Hebrew and Modern Hebrew on the one hand, Classical Arabic and Egyptian Arabic on the other. Ancient Hebrew was classified as VS while Modern Hebrew was classified as SV by the trained classifier. Classical Arabic was classified as VS while Egyptian Arabic was classified as SV. This indicates that the relative lengths of nouns and verbs, when frequency is taken into account, can predict word order from corpora, with various implications for our understanding of language processing and language change. This is a striking finding, and I am unaware of any other research that has been able to show predictive validity of a model related to word order change. However, it is possible that this result is conditioned to some degree by the fact that these are Semitic languages spoken in the same geographical area, so we now turn to an investigation of whether descent or location provides a better explanation for the correlation of noun and verb lengths with word order.
2.4 Testing variance predicted by word length
Dunn et al. (2011) showed that descent from a common ancestor was a better predictor of many grammatical patterns than other explanations. Although Jäger & Wahle (2021) counter this finding with additional data, we can also assess this claim using the current dataset to see if descent is more explanatory of word order for a given language than the lengths of nouns and verbs. Additionally, in more recent work, Naranjo & Becker (2022) show that areal affects are as important as family membership in predicting linguistic phenomena, finding that the results in Dunn et al can be just as easily predicted by language area. This is also a claim that we can test with our dataset. Accordingly, this investigation can proceed with several (motivated) simplifying observations.
First, for each language we can take the highest level node of family membership that is known based on current scholarship (via Glottolog). [^17] Glottolog (Hammarström et al. 2024; https://www.glottolog.org) contains expert annotations for language family affiliations for over 7,000 languages, along with additional aggregated data from a variety of sources. This is because all modern languages are the same age, and while their pathways/histories may differ, this does not impact whether they can be considered part of the same phylogenetic tree. That is, if descent is more explanatory than word class length, we would expect languages belonging to the same tree to have a higher likelihood of sharing the same word order than not. Second, we can use the macroareas identified by Glottolog as well as the actual GPS coordinates of languages to control for areal affects. Third, we can run various analyses with differing sets of data in order to see how this affects the statistical robustness of the findings. It may be, for example, that using a smaller sample gives a different result than a larger or more diverse one. Here, the more languages we have for a language family or geographical location, the higher the likelihood that any relationship identified is not due to chance. Finally, we can remove languages with “free” word order from consideration, as this leaves us with a binary distinction in word order (SV, VS) that opens up more possibilities for statistical tools.
In particular, a binary dependent variable for word order means we can use hierarchical linear regression, which allows us to observe the relative impact of particular features that are theoretically relevant to the question at hand by seeing how they influence the performance of successive models. In the case of word order, we can train a regression model with a feature known to be relevant to word order (the N1 ratio), and then subsequent models that incorporate additional features of interest (“language area/location”, “family membership”, “noun/verb lengths”). We can then compare the models to see whether the additional features increase the variance explained in the data via the F-value: an increase in F-value (change) and the significance of this change indicates more/additional variance explained by the particular feature added in a model.
Accordingly, for all languages in the tagged PBC I retrieved the family membership and geographical information from Glottolog, filling in missing data via Ethnologue and other sources. Then I filtered the data on four conditions: a) languages found in Dunn et al that are also present in the tagged PBC (106 languages), b) languages found in the tagged PBC that are in the four families investigated by Dunn et al (539 languages), c) languages in families with more than 75 members in the tagged PBC (962 languages), and d) languages in families with 2 or more members in the tagged PBC (1,486 languages). For each of these conditions, I ran a hierarchical linear regression, adding features of interest progressively to see how they impacted the fit of the successive models and whether particular features account for additional variance.
The base model (Model 1) uses the N1 ratio as the primary feature for classification. This is because we know from comparison with expert determinations of word order (see footnote 5 above) that the N1 ratio significantly correlates with the classification of languages as SV/VS. Due to the strong relationship observed between language area and typological properties of language (see Naranjo & Becker 2022), Model 2 then includes GPS coordinates and macroarea identifiers for each language as features. Model 3 additionally incorporates descent (language family membership) as a feature (as per Dunn et al. 2011), and Model 4 includes the average lengths of nouns and verbs (as per the relationship identified in this paper). If descent is more explanatory than the lengths of nouns/verbs, this should be clear from the differences between the respective models for each condition. Results are displayed in Table 2.
a) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
1 (N1 ratio) 106 45.85 7.8e-10
2 (Lg area) 106 20.40 2.4e-12 8.58 3.9e-05**
3 (Family) 106 16.70 5.8e-12 1.48 0.227
4 (N/V len) 106 13.40 4.9e-12 3.27 0.042*
b) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
1 (N1 ratio) 539 230.41 1.4e-43
2 (Lg area) 539 100.00 1.7e-63 39.86 2.9e-23**
3 (Family) 539 80.71 5.3e-63 2.46 0.118
4 (N/V len) 539 64.75 3.5e-67 14.57 6.9e-07**
c) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
1 (N1 ratio) 962 426.74 9.9e-79
2 (Lg area) 962 220.49 4.2e-134 105.35 6.1e-59**
3 (Family) 962 176.22 6.6e-133 0.04 0.839
4 (N/V len) 962 130.93 8.3e-135 9.69 6.7e-05**
d) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
1 (N1 ratio) 1486 586.46 1.8e-109
2 (Lg area) 1486 193.61 1.3e-133 45.19 6.0e-28**
3 (Family) 1486 155.10 1.2e-132 1.03 0.309
4 (N/V len) 1486 115.43 3.5e-135 11.02 1.7e-05**
Table 2: Hierarchical Linear Regressions under 4 conditions
From these analyses we can see that, as expected, the N1 ratio significantly predicts word order under all conditions. Additionally, language area is a significant predictor of word order that adds to the strength of the classification. Language family (descent) does not add additional predictive power to our model under any conditions (over language area), whereas the length of nouns and verbs does. This analysis shows that the length of nouns and verbs in a language accounts for more variance and at the same time is more significant than that language’s family membership in determining its word order classification, under all test conditions. When we remove language area from the linear regression this is still the case, [^18] See additional statistical results at the linked repository (Fn 4). indicating that while language area is a stronger predictor than family membership, lengths of nouns and verbs is yet a stronger one.
This largely aligns with Jäger and Wahl’s (2021) finding regarding the (lower) correlation between descent and linguistic structures in comparison to other explanations, with the caveat that Jäger and Wahl included language isolates, which is likely to greatly diminish the affect of descent. If a large number of languages in a sample are not related to others, clearly descent can have only a minimal impact on word order. By including only language families with two or more members (which excludes 111 languages in the taggedPBC), I show that descent does not explain much variance above and beyond language location/area, in contrast with noun and verb lengths. This also suggests that language family/descent is highly correlated with linguistic area, which conforms with expectations related to human history and migration. Additionally, the amount of variance explained seems to depend on the number of languages (and language families) in the sample, highlighting the need for more data for languages from under-represented language families.
3 Discussion
That arguments overall are shorter in languages of the tagged PBC, across all word orders, suggests that there are certain (non-Noun) arguments affecting our observations. Particularly, the effect of PRON seems quite likely, whereby a tendency for pronouns to be shorter crosslinguistically, along with a preference for pronouns as arguments, may impact the observation that arguments are shorter overall than predicates. This pattern aligns with the processing claim that ‘heavier’ words (or word classes in this case) occur later in a sentence, partly validating Hypothesis 1. The counter-argument, also from the data, is that nouns are longer than verbs in SV languages and shorter than verbs in VS languages when frequency is considered, validating Hypothesis 2. This suggests that speakers prefer to place heavier (more “contentful”, or perhaps “novel”) constituents before lighter ones crosslinguistically in many cases.
These seemingly antithetical results partially validate both the “weight”-based theory and the “surprisal” theory, highlighting the interaction of two pressures: one whereby speakers prioritize lighter constituents (possibly for the purposes of production/processing efficiency), and one where they prioritize heavier constituents (possibly for informational purposes). The two perspectives may in fact be looking at the same data from different angles. On the one hand, grouping all arguments into a single class highlights the overall impact of shorter strings, since the most frequently used arguments (pronouns, demonstratives) are typically shorter crosslinguistically. On the other, more “content-ful” word types that aid planning and inference or predictability (nouns, verbs) will typically be longer crosslinguistically.
These two pressures can be (and, in fact, have been) considered as two different kinds of “efficiency”: production efficiency vs communicative efficiency. Under this view, production efficiency is related to length of words (as per Wasow 2002, 2022; Hawkins 2014) - speakers use shorter words whenever they can to efficiently produce utterances (see also Mahowald et al. 2013). The counter-pressure, communicative efficiency, is related to making a speaker’s utterances more readily accessible/interpretable by recipients (as per Levshina & Moran 2021; Levshina 2023) - speakers prioritize words that contain more meaningful, informative content in order to aid communication. Once context has been established, these longer words can be replaced by more efficient (shorter) terms such as demonstratives or pronouns. [^19] For example, there is a common crosslinguistic pattern of using a noun first to establish reference, then switching to pronouns, i.e. "that tree, it is big, it will fall…" See also discussion of the (lack of) encoding of subjects crosslinguistically in Berdicevskis et al. (2020).
I would argue, however, that conceptualizing “efficiency” in these two distinct ways, as currently done in the literature, runs the risk of confusing the two. It is in fact not efficient (from a processing perspective) to maximize information, as this also tends to maximize complexity. Instead, I propose a different framing of these principles.
3.1 A “Min-Max” theory of processing
The current proposal re-formulates or refines Hawkins’ PGCH to incorporate (increased) complexity as a motivation for the order of constituents, bearing some resemblance to proposals in Levshina (2023). I frame this as (to borrow a term from computer gaming) the “Min-Max” theory of processing (schematized in Figure 3). The idea here is that speakers seek to minimize their processing effort while maximizing the information in their utterances. These are competing/parallel pressures that speakers are constantly balancing in production (or “performance”) in order to choose the best strategy for their immediate goals. As speakers produce variants along a cline of grammaticality, these variants compete on the basis of multiple factors, including how well such variants support Min-Max-ing in relation to other variants, leading to language change.
Figure 3: Min-Max schema
These two competing pressures offer a more complete (and parsimonious) explanation for word order change (and possibly other kinds of linguistic change) than existing theories of language. In one possible scenario the frequent use of short pronominal forms encourages drift toward an argument-initial order, and nouns subsequently lengthen as they provide more information to speakers and move toward an initial position. This, combined with the preponderance of shorter pronominal forms crosslinguistically, may provide an explanation for why the majority of languages reported in the world are SV. Conversely, if a language has or develops longer verbs, the preference for these heavier constituents to occur first could put pressure on the language to maintain or develop predicate-initial order, which would explain why some VS languages are quite stable. In each case the direction of development would depend on a number of factors that have been implicated in discussions of language change and grammaticalization (see Heine & Kuteva 2002; Matras 2011; Narrog & van der Auwera 2011). However, the effect of word length is clearly a major part of the explanation for why we can observe (and predict) such a difference in word order between Ancient and Modern Hebrew as well as Classical and Egyptian Arabic via their respective corpora.
The existence of multiple competing pressures affecting the development of language, as described by this Min-Max theory, is supported by work in clinical psychology (e.g. Rezaii et al. 2023) and language modeling (e.g. Hahn & Xu 2022), as well as findings from information theory (Zaslavsky 2020; Tucker et al. 2025). In a study of patients with nonfluent aphasia, Rezaii et al. (2023) showed that patients with this condition compensate for their deficit in producing complex grammatical structures by increasing the information content of their utterances, dropping high-frequency (function) words in favor of low-frequency (content) words. By modeling variations in grammatical structure based on corpora from 80 languages, Hahn & Xu (2022) found that the resulting models were less optimized for dependency length and information locality than the original structures of each language, indicating that languages evolve to efficiently balance the two considerations. [^20] Notably, Hahn & Xu propose a theory that the interaction between Dependency Length (DL) and Information Locality (IL) accounts for word order variation and change, but neither of the two concepts, as expressed, make claims about directionality/order, only about the relationships between (internal) constituents. In contrast, the Min-Max theory makes the claim that speakers balance information content (as reflected by longer nouns in SV languages and longer verbs in VS languages) with processing/production concerns, and that this interaction is what drives the evolution of word order. As nouns lengthen or verbs shorten in a language, this will affect their placement in a stream, which then impacts the overall order observed in surface structures. That is, the Min-Max theory makes a universal claim regarding language-external pressures, while DL and IL are language-internal concepts. These and other studies indicate that language users are sensitive to efficiency as it pertains both to processing and information content.
The sensitivity of language users to processing efficiency and information content has also been a major topic of investigation in information theory, which is concerned in part with the encoding and decoding of information streams. One information-theoretic proposal is that semantic systems evolve following an “Information Bottleneck” principle (IB; Tishby et al. 1999), whereby “optimal representations satisfy a tradeoff between compressing X, i.e. minimizing the representational complexity, and maximizing the relevant information about Y” (Zaslavsky 2020: 14). Work by Zaslavsky shows that IB can account for the development of color naming systems in language among other kinds of sytems, and additional research shows the utility of IB in machine learning and distributional models of semantics (Slonim & Tishby 2000).
The implications of this principle have most recently been examined in agent-based emergent learning simulations with the addition of a “utility” principle (Tucker et al. 2025) – the interaction between “utility”, “informativeness”, and “complexity” allows more human-like communication to emerge among agent models in their study. Such a finding is very much supported by the results reported here, though I would argue that the Min-Max theory offers a more parsimonious account of human behavior as related to language specifically. While the current corpus evidence applies at the word level in relation to basic word order, it seems likely that the minimization of processing difficulty would interact with the maximization of information content at multiple (hierarchical) levels. Speakers min-max in order to select words in a stream (“informativeness” and “complexity” as per Tucker et al.), while simultaneously min-maxing to select relevant structures related to their communicative goals (Tucker et al.’s “utility”).
The Min-Max theory also bears some resemblance to Levshina’s (2023) proposals, described in terms of communicative efficiency and accessibility. However, such indicators cannot fully explain certain patterns in grammar such as “harmonic” word order. A recent proposed explanation for such patterns involves frequency of word classes, whereby Mansfield & Krapp (2025) suggest that “harmonic ordering is really a frequency effect: as replication converges on a phrase structure with a consistent linear order, the most frequent word class tends to be at one edge.” Although they did not test token length in their model, it is notable that in the current study frequency is a major factor in identifying the relationship between length and word order, suggesting that the most frequent word class in their model may also be the longest element of the domain they are testing, which would be driven by a maximization of information principle. The Min-Max theory also seems to provide an explanation for non-parallel (hierarchical) harmonic word order, which is not explained by their model.
The Min-Max theory may additionally provide an explanation for why word forms remain long (or lengthen), as described for particular languages (see Cristofaro 2021; Haspelmath 2021; Petré 2017). Here, again, the maximization of information principle would support the development or retention of particular (long) word forms, as speakers balance this need against the minimization of processing difficulty. As noted above, this proposal is supported by evidence from color-naming systems, which show patterns in keeping with the Information Bottleneck principle. The Min-Max theory would suggest that this principle does not only hold for “forms” (of, i.e. color words), but also governs “behavior”: speakers behave in ways that are consistent with maximizing information in streams (via word order), while minimizing processing difficulty.
As noted in section 2, such a theory can be conceived of as a universal (underlying) pressure that applies generally to how language is realized. From an innate perspective, speaker choices impacted by such pressures would lead to a re-analysis or re-formation of rules that guide surface realizations. From a usage-based perspective such pressures would lead to the development of new surface variants. These are not necessarily opposing views (see Rastelli 2025), and the outcomes are quite similarly dependent on language being viewed as a complex adaptive system, whereby diachronic changes may have a combination of both “result-oriented” and “source-oriented” explanations (see Cristofaro 2019; Schmidtke-Bode & Grossman 2019).
Another observation from the results reported here is that information content seems to trump length. We find that long nouns are used with more frequency in SV languages, while long verbs are used with more frequency in VS languages. This indicates that (maximizing) information content is more important to speakers than (minimizing) processing complexity. This could explain the “agent/subject-first preference” in experiments (Sauppe et al. 2023; Culbertson et al. 2012), as the “information maximization” pressure wins out in cases where this most efficiently meets communicative goals.
Further, the data shows that all languages have a relatively high proportion of sentences in which the argument occurs before the predicate. In VS languages, the proportion of argument-initial to predicate-initial sentences is 1/2 or more (min: 0.56 n/v, 0.65 a/p), highlighting that even VS languages still place the noun/argument first in many constructions. Similarly, while all SV languages have a high proportion of N1 sentences (max 2.43 n/v, 7.17 a/p), the number of V1 constructions is non-zero, supporting a gradient approach to word order (Levshina et al. 2023).
The fact that the length of particular word classes can predict the basic word order of a language is a novel finding, leading to some general predictions: if, on average, nouns are longer than verbs in a language, it is highly likely to have SV word order. If, on the other hand, verbs are longer, it is highly likely to have VS word order. The effect of frequency cannot be ignored either, highlighting the importance of conducting large-scale corpus based studies to investigate this question. Since languages are used by speakers and such speakers use certain words more than others, we can only observe this frequency effect (or “weight”) of word class length via corpora.
The idea that speakers balance multiple considerations in production is not new (Pike 1967; Jackendoff & Pinker 2005; Seržant & Moroz 2022; Hörberg & Sjons 2023) but what is novel is the idea that all humans experience a specific universal pressure to balance production/processing efficiency with information content. This pressure is observable crosslinguistically, and would apply at multiple processing levels, regardless of whether a speaker is accessing a universal grammar (“innate”) or inferring structure on the basis of the input they receive (“usage”). Following on from Zipf’s and subsequent observations, there is no doubt that the length of strings is an important factor in processing streams, but clearly not all strings are created or treated equally by speakers. This study shows that nouns are treated differently (or at least have a different distributional relationship) than other kinds of arguments, which leads to different realizations of word order crosslinguistically. The pressures identified here have wide-ranging implications and applications for any conceptualization of how language operates, whether in terms of understanding the intricacies of individual linguistic scenarios, designing intelligent systems that use language, or developing clinical interventions to assist speakers in various ways.
3.2 Limitations
One limitation in this study is the use of automated processes to identify part of speech tags in the data. Ideally, all the data should be hand-annotated for parts of speech, which would allow for better comparison of the different word categories across languages. Given the time consuming nature of this task and the need for expert annotators, this is not yet possible, but ongoing research could address this by sourcing experts to annotate the data. Although I have shown (in Ring 2025 and via the linked repository) that the translation/alignment method does correspond well to hand-tagged data, it may be that this could be improved by tagging a small amount of data for each language in the PBC in order to train POS taggers that perform better than the current translation method.
Another limitation is that the dataset used, while exceedingly large in comparison to previous efforts, still only represents about a quarter of the world’s languages. Since the majority of the world’s languages are not represented here, it is not clear that these findings can be fully generalized to all languages of the world, though I have shown in section 2.4 that much can be learned from sampling. Additionally, the contrast between intransitive SV/VS/free word order is somewhat of an oversimplification of word order patterns. Given the current obstacles to annotation, it is a necessary simplification for this study, however future work could expand to a more diverse set of word/constituent orders and potential relationships with word length and/or complexity. Given the nature of the dataset used, it is also possible to pursue a gradient approach to word order, as advocated by Levshina et al. (2023). The issue of word complexity also could use further expansion to cover a broader range of possibilities than simply the length of a word based on a set of (roman) characters.
A final note is that our conclusions are only as good as the data that we have available to us. The tagged PBC is a significant step in the right direction, toward a representative sample of large tagged corpora for languages of the world. Systematic investigation of this dataset can give us insight into language typology (the current state of language), as well as pressures that impact this realization in diachrony, both in terms of how languages have developed (their history) and how they might develop (their future). On the basis of this investigation I have proposed a parsimonious “Min-Max” theory of processing and information structure that underlies language evolution, whereby speakers minimize processing effort while maximizing information gain. This proposal is in keeping with at least one existing information-theoretic proposal (IB) as well as theories of processing (PGCH; surprisal) and can explain the correlation between the length of nouns/verbs and basic word order, given that these features allow us to predict the word order of a language. It is as yet unclear whether such a principle would apply at multiple hierarchical levels within language, which is an area that deserves further careful investigation with datasets such as the tagged PBC. For now, I have demonstrated that noun/verb lengths are more explanatory than descent or geography and that for languages with corpus data, word length predicts word order.

### v2

@@
 Abstract
-Current theories of language propose an innate (Baker 2001; Chomsky 1981) or a functional (Greenberg 1963; Dryer 2007; Hawkins 2014) origin for the surface structures (i.e.
-word order) that we observe in languages of the world, while evolutionary modeling (Dunn et al.
-2011) suggests that descent is the primary factor influencing such patterns.
-Although there are hypotheses for word order change from both innate and usage-based perspectives for specific languages and families, there are key disagreements between the two major proposals for mechanisms that drive the evolution of language more broadly (Wasow 2002; Levy 2008).
-This paper proposes a universal underlying mechanism for word order change based on a large tagged parallel dataset of over 1,500 languages representing 133 language families and 111 isolates.
-Results indicate that word class length is significantly correlated with word order crosslinguistically, but not in a straightforward manner, partially supporting opposing theories of processing, while at the same time predicting historical word order change in two different phylogenetic lines and explaining more variance than descent or language area in regression models.
-Such findings suggest an integrated “Min-Max” theory of language evolution driven by competing pressures of processing and information structure, aligning with recent efficiency-oriented (Levshina 2023) and information-theoretic proposals (Zaslavsky 2020; Tucker et al.
+A fundamental concern in linguistics has been to understand how languages change, such as in relation to word order.
+Since the order of words in a sentence (i.e.
+the relative placement of Subject, Object, and Verb) is readily identifiable in most languages, this has been a productive field of study for decades (see Greenberg 1963; Dryer 2007; Hawkins 2014).
+However, a language’s word order can change over time, with competing explanations for such changes (Carnie & Guilfoyle 2000; Crisma & Longobardi 2009; Martins & Cardoso 2018; Dunn et al.
+2011; Jäger & Wahle 2021).
+This paper proposes a general universal explanation for word order change based on a theory of communicative interaction (the Min-Max theory of language behavior) in which agents seek to minimize effort while maximizing information.
+Such an account unifies opposing findings from language processing (Piantadosi et al.
+2011; Wasow 2022; Levy 2008) that make different predictions about how word order should be realized crosslinguistically.
+The marriage of both “efficiency” and “surprisal” approaches under the Min-Max theory is justified with evidence from a massive dataset of 1,942 language corpora tagged for parts of speech (Ring 2025), in which average lengths of particular word classes correlates with word order, allowing for prediction of basic word order from diverse corpora.
+The general universal pressure of word class length in corpora is shown to give a stronger explanation for word order realization than either genealogical or areal factors, highlighting the importance of language corpora for investigating such questions.
+1 Introduction
+There are two major approaches to studying word order.
+The first (typological) approach is to observe word orders in natural languages and identify patterns whereby the orders of certain constituents are correlated with other grammatical patterns.
+This approach has developed multiple crosslinguistic databases with expert determinations of word order (WALS, Dryer & Haspelmath 2013; AUTOTYP, Bickel et al.
+2023; Grambank, Skirgård et al.
+2023) and has identified statistical universals in which the orders of certain elements increase the likelihood for other patterns to be present in a language (Greenberg 1963; Dryer 2007, 1992, 2011; Verkerk et al.
 2025).
-1 Introduction
-Theories of language should describe not only the current state of languages, but also explain how they have developed (their history) and predict how they will change (their future).
-To date, however, existing theories have primarily focused on describing the present state of language, proposing that the current diversity in surface structures or patterns that we observe in languages of the world, such as for word order, have either an innate (Baker 2001; Chomsky 1981) or functional (Greenberg 1963; Dryer 2007; Hawkins 2014) source.
-More recent diachronic work has tried to answer the second question on how languages develop.
-From an “innateness” perspective, surface word orders are seen to be derived from an underlying (universally shared) structure, whereby parameters for surface realizations change over time (Carnie & Guilfoyle 2000; Crisma & Longobardi 2009; Martins & Cardoso 2018), while from a “functional” perspective surface word orders are the result of cumulative changes in linked structures (sometimes called ‘constructions’) that compete for adoption in individual languages (see Gerritsen & Stein 1992; Leino & Kuningas 2006; Barðdal & Gildea 2016).
-Importantly, none of these perspectives are particularly explanatory of drivers in language evolution that affect all languages similarly.
-The closest observations in this regard come from evolutionary modeling, where descent from a common ancestor was found to be the most explanatory factor in predicting word order (Dunn et al.
-2011; though see Jäger & Wahle 2021; Naranjo & Becker 2022 for alternate perspectives), and language processing, where word length (Piantadosi et al.
-2011; Wasow 2022) and surprisal (Levy 2008) have been shown to impact the realization of word order.
-There are also more general claims that “efficiency” (variously defined) explains grammatical (Hawkins 2014; Levshina 2023) or semantic (Zaslavsky 2020) change.
-However, the fact that languages tend to display word orders like their family members does not explain the variation in word orders that we find within families, and perspectives on processing make very different predictions: the literature on word length predicts that short/light words will occur first in a sentence, while the literature on surprisal predicts that long/heavy words will occur first.
-Further, the literature on efficiency does not make actual testable cross-linguistic predictions about basic word order.
-There are also other issues to consider, based on observation of the languages for which we have data.
-Typological surveys (Dryer & Haspelmath 2013; Skirgård et al.
-2023; Bickel et al.
-2023) reveal that over 80% of the sampled languages have a Subject-initial basic word order, while only about 10% have a Verb-initial order.
-Further, there are various correlations to be found between the orders of clausal elements, such that basic word order is related to how other clause types are ordered.
-While all languages allow variation in word order to different degrees, the word order most often found in declarative sentences can predict (to some degree) the ordering of other grammatical elements in a language (Dryer 1992, 2011).
-It is not completely clear why one basic constituent order correlates better with one particular order of (e.g.) noun phrase elements than another, nor why some basic word orders are less well-represented crosslinguistically.
-This raises the question whether there are underlying processes that drive these observed tendencies.
-Some explanation clearly comes from descent, i.e.
-that languages tend to resemble their ancestral forms.
-However, the word order changes that languages undergo over time suggests other pressures in terms of how complex linguistic elements are and how these constituents are processed in the brain (Maurits & Griffiths 2014; Schouwstra & de Swart 2014, see also Piantadosi et al.
-2011; Bickel et al.
-2015; Ding et al.
-2017; Fernandez 2016; Jaeger & Tily 2011; Tily 2010; Sinnemäki 2014).
-Answers to these questions have been elusive in part due to a lack of data that would facilitate crosslinguistic investigations.
-This is because such investigations require data that is annotated for various word types or dependencies, which is unavailable for the majority of the world’s languages.
-The largest sample of languages used to investigate these questions looked at annotated corpora in 80 languages (Hahn & Xu 2022), the majority of which are from the Indo-European language family.
-Such data risks biasing findings toward a particular group of languages (Blasi et al.
+The second (processing) approach focuses on experimental evidence for how constituents in particular languages are processed in the brain.
+This approach has identified potential universal tendencies or preferences in processing, notably a preference for “agents first” (Culbertson et al.
+2012; Futrell et al.
+2015a; Sauppe et al.
+2023), “efficiency” (Zipf, 1949; Wasow, 2002; Futrell et al., 2015b), and “surprisal” (Hale 2001; Levy 2008).
+While the typological observations do not make predictions about a language’s basic word order (the universals take basic word orders as a starting point and make further predictions, see Verkerk et al.
+2025), the processing research does imply some particular claims and predictions about how words should be ordered in languages of the world.
+For example, the “agents first” (or “Subject first”) bias found in experiments suggests that there is a universal psychological pressure for speakers of all languages to produce agent arguments (“do-ers” of actions, typically represented by animate nouns and pronouns) before other kinds of constituents in sentences.
+This aligns with observations of the dominant crosslinguistic word order being Verb-final or Subject-initial (Schouwstra & de Swart 2014; Schouwstra et al.
+2022).
+At the same time, findings from language processing studies suggest that other pressures mediate the realization of word order patterns, namely in relation to “efficiency” and “surprisal”.
+The efficiency literature largely follows from Zipf’s (1932; 1949) observations that a) languages show a logarithmic distribution whereby the most frequently used words occur much more often than the next most frequent, and that b) the most frequent words in a language are also the shortest.
+The observation that shorter (or less complex) elements are more likely to occur initially in a sentence, while longer (or more complex) elements occur later, can be traced to Behaghel (1909), and the general idea is that this supports efficient communication.
+In contrast to this view, the surprisal literature focuses on how salient or predictable items are, suggesting that longer (or more complex) words are more likely to occur first in a sentence since they improve predictability of following constituents (Hale 2001; Levy 2008; Sauppe et al.
+2021).
+This is supported by clinical evidence (Rezaii et al.
+2023) and information theoretic studies (Zaslavsky 2020; Tucker et al.
+2025).
+Such studies have implications for particular languages regarding the ordering of constituents.
+For example, if a language has longer verbs than nouns, we would expect the language to be Verb-final based on the efficiency literature (longer or more complex words last), whereas the surprisal literature would predict it to be Verb-initial (longer or more complex words first).
+These competing predictions represent hypotheses that have yet to be tested to any great degree (though some attempts have been made, at least in relation to dependency structures, see Jing et al.
+2021; Hahn & Xu 2022).
+The lack of testing highlights the complexity of investigating pressures that influence the evolution and realization of language structures, as well as the dearth of datasets and databases that could address these questions.
+[^1] Datasets that have been used by previous research include the Universal Dependencies Treebanks (Zeman et al.
+2024) and ODIN (Xia et al.
+2016) projects.
+The UDT contains 160+ hand-annotated languages available for crosslinguistic investigation, but these are not parallel texts and the number of sentences in each language is variable, with some having fewer than 200 sentences in the v2.14 release.
+The ODIN project (and its re-packaged form via LECS: https://huggingface.co/datasets/lecslab/glosslm-corpus) has interlinear glossed text (IGT) from over 1,800 languages, but the majority of these are single-word/line examples, and only 300 of the languages in their dataset have more than 100 sentences.
+Additional cross-linguistic investigations using tagged corpora include Hahn et al.
+(2020: 51 languages), Jing et al.
+(2021: 71 languages), and Hahn et al.
+(2022: 80 languages).
+For each of these studies the majority of languages are from the Indo-European language family.
+Besides limiting generalizability of findings, such data risks bias toward a particular group of languages (Blasi et al.
 2022), in much the same way that other scientific findings have historically been biased toward WEIRD demographics (Henrich et al.
-2010), leading to recent calls for more expansive datasets to examine linguistic questions (Benítez-Burraco 2025; Levshina et al.
+2010).
+In particular, questions regarding the length of constituents in relation to their order in sentences requires a large number of sentences tagged for parts of speech (at minimum).
+Given that languages develop over time via cultural evolution, it is also difficult to separate general cognitive pressures from historical (genealogical) or areal factors.
+This means that inferring crosslinguistic or universal principles necessitates both a large number of languages as well as sufficient diversity of languages to establish statistical validity.
+Additionally, the fact that languages develop over time through communicative interaction requires that any explanation for universal principles be related to such interactions.
+That is, in an evolutionary linguistic paradigm it is individuals (speakers) in a population (language) that drive variation and select particular linguistic variants for propagation.
+The selection process itself is likely motivated by various factors, two of which are combined in the Min-Max theory of language behavior.
+The Min-Max theory proposes that in production speakers select variants based on two competing pressures: the pressure to minimize effort and the pressure to maximize information (Figure 1).
+This language behavior is considered to be core to what drives selection and propagation of linguistic constructions, and thus change over time (in, i.e.
+word order).
+Figure 1: Min-Max schema
+The current paper takes advantage of a massive tagged parallel corpus to show how these pressures are realized in actual language data in relation to word order.
+While the dataset is still under development, the baseline data is tagged for parts of speech, which allows us to derive a proxy measure for word order from individual language corpora, as well as to measure other linguistic properties such as the average length of various word types.
+Results show that basic word order in intransitive clauses is highly correlated with the length of particular word classes, allowing for the prediction of a language’s word order based on corpora not in the dataset and explaining variance better than both genealogy and areal factors.
+In the remainder of this paper I first highlight the properties of the dataset (§2).
+I then discuss three studies that investigate word length in relation to word order (§2.1), the predictive validity of word lengths for word order (§2.2), and the variance predicted by word length in regression models (§2.3).
+I conclude with a discussion of the implications of these studies in relation to language evolution (§3), presenting the Min-Max theory of language behavior (§3.1) as well as some limitations and future directions regarding the current effort (§3.2).
+2 Methodology and hypotheses
+The taggedPBC (Ring 2025) [^2] https://github.com/lingdoc/taggedPBC - see fn 5 for a link to a minimal OSF repository that allows for replication of the studies in this paper.
+contains pos-tagged parallel text data in 1,942 languages (roughly 25% of the world’s languages), with over 1,800 parallel verses for more than 1,850 of those languages.
+The dataset represents 155 language families and 78 isolates, dwarfing previously available resources, and the tagging methodology (via crosslingual tag transfer) has been shown to correspond well with existing automated taggers for high-resource languages (as found in the SpaCy and TranKit NLP libraries).
+[^3] https://spacy.io ; https://nlp.uoregon.edu/trankit Additionally, a novel measure extracted from individual corpora in this dataset (the “N1 ratio”) is significantly correlated (p < 0.001) with expert determinations of a language’s basic word order for intransitive clauses (“SV”, “VS”, “free”; Figure 2) as identified in the typological databases WALS (Dryer & Haspelmath 2013), Grambank (Skirgård et al.
+2023), and Autotyp (Bickel et al.
 2023).
-The current study addresses this by developing a large corpus of texts tagged for word class.
-With data sourced from the Parallel Bible Corpus (PBC; Mayer & Cysouw 2014), machine learning techniques allow for automatic tagging [^1] Automated part of speech tagging was validated via comparison with existing taggers for high-resource languages (SpaCy, 16 languages - avg nouns: 0.83, verbs: 0.74; Trankit, 31 languages - avg nouns: 0.83, verbs: 0.76) and gold-standard tagged data in the Universal Dependencies Treebanks (UDT; Zeman et al.
-2024; 106 languages; https://universaldependencies.org).
-See Ring (2025) for more details.
-of word classes in 1,597 languages (roughly 20% of the world’s languages) that represent 133 families and 111 isolates.222‘Isolates’ being defined here as languages in the corpus that share no known family membership with other languages in the corpus.
-Individual corpora in the majority of these languages (over 1,500) contain more than 1,800 parallel sentences, and the smallest corpus contains more than 700 sentences.
-[^3] For clarity, the UDT has 160+ hand-annotated languages available for crosslinguistic investigation, but these are not parallel texts and the number of sentences in each language is variable, with some having fewer than 200 sentences in the v2.14 release.
-The ODIN project (and its re-packaged form via LECS: https://huggingface.co/datasets/lecslab/glosslm-corpus) has interlinear glossed text (IGT) from over 1,800 languages, but the majority of these are single-word/line examples, and only 300 of the languages in their dataset have more than 100 sentences.
-Additional cross-linguistic investigations using tagged corpora include Jing et al.
-(2021: 71 languages) and Hahn et al.
-(2020: 51 languages).
-This dwarfs both the linguistic data and the language diversity available from other sources and means that crosslinguistic investigations using this dataset can be confident that they are comparing similar contexts.
-The dataset (the tagged PBC) and explanations regarding its development, along with Python code for replicating the results in this paper, can be found at the linked repository [^4] https://github.com/lingdoc/taggedPBC and is described at greater length in Ring (2025).
-This dataset can be used to investigate the crosslinguistic questions highlighted above regarding whether there is some pressure (beyond descent) that explains the evolution of word order in languages of the world.
-If there is indeed a mechanism that underlies the observed surface structures in present-day languages, this should also be one that reveals historical processes and allows us to predict future developments.
-Specifically, by deriving a novel corpus-based measure of word order (the “N1 ratio”) [^5] The N1 ratio was validated via comparison with existing classifications of languages in the tagged PBC and three typological databases: the World Atlas of Language Structures (WALS; Dryer & Haspelmath 2013; https://wals.info/), the AUTOTYP Database (Bickel et al.
-2023; https://github.com/autotyp/autotyp-data), and Grambank (Skirgård et al.
-2023; https://grambank.clld.org/) where the N1 ratio significantly differentiated (p ¡ 0.001) between SV/VS/free classifications in all three databases.
-See Ring (2025) for more details.
-from this dataset, the current paper investigates the relationship between basic word order and the lengths of particular word classes (nouns/arguments and verbs/predicates) crosslinguistically.
-Results show that the lengths of these word classes is significantly related to word order, that they can predict historical word order change in two different lineages, and that they account for more variance than family membership (descent) or areal patterns (geography) when examined using hierarchical linear regressions.
-Such findings suggest the possibility for a unified (“Min-Max”) theory of language evolution that can explain both observed surface structures in languages of the world as well as how such patterns have developed diachronically.
-2 Complexity, processing, and word order
-Linguistic complexity can take many forms, with many considerations regarding its measurement (see Jaeger & Tily 2011). As noted by Joseph and Newmayer (2012: 341):
-“how linguistic complexity is conceived of, and how it might be identified and measured, has itself never been agreed upon by linguists.
-It cannot be neatly separated from theoretical considerations of where language structure ends and language processing (or use) begins, or where variation (whether social or stylistic) is located.”
-Some current efforts to measure linguistic complexity focus on quantifying the overall complexity of a language (Bentz et al., 2016, 2017; Baerman et al., 2015, 2017), and such measures may be under considerable debate (see Bentz et al., 2023).
-Additional research suggests that word complexity can be measured by word length (see Lewis & Frank, 2016; and see Piantadosi et al., 2011; Meylan & Griffiths, 2021; Koplenig et al., 2022 for further discussion in relation to frequency in corpora).
-Words which are formed from longer sequences of phonemes are more difficult to process, and this is also related to their conceptual complexity.
-The fact that word length is related to a word’s complexity is important for our understanding of word order, i.e.
-how closely is the ordering of constituents related to how we process words/structures of varying lengths?
-Research into how speakers process words has identified several constraints and tendencies that have various implications for how languages develop.
-Hawkins (2014) summarizes this literature in arguing for a comprehensive theory (the Performance-Grammar Correspondence Hypothesis; PGCH) based firmly in the idea of efficiency as an explanation for a large number of observed patterns and historical changes cross-linguistically.
-[^6] Levshina (2023) argues for another explanation termed “communicative” efficiency, which identifies “accessibility” as a factor in language use.
-This seems to overlap to some degree with the effects identified by “surprisal”.
-As Levshina’s observations do not make specific predictions about word order crosslinguistically, I revisit them in the discussion.
-His essential claim is that speakers minimize complexity (in forms, distance) in order to speed up processing, a claim that aligns with observations about constituent “weight” (Wasow, 2002; Zipf, 1949) and “dependency length minimization” (Futrell et al., 2015).
-However, this claim contrasts with findings from Surprisal Theory (Hale 2001; Levy 2008) that speakers maximize complexity, i.e.
-that more complex forms are preferred by speakers since they support information content/retrieval and reduce entropy in utterances (see Rezaii et al.
-2023).
-These two opposing perspectives on processing (the minimization vs maximization of complexity) make two different predictions for word order crosslinguistically.
-The PGCH claim (and Wasow’s) is that “heavier” or more complex constituents are typically “put off” until later in the sentence, while less complex constituents are put first.
-This predicts that a language in which nouns in general are easier to process than verbs should display SV order, while one in which verbs are easier to process should display VS order.
-Yet evidence from languages like Japanese, which places heavier phrases closer to the beginning of the sentence (Hawkins, 1994; Yamashita & Chang, 2001; Yamashita, 2002; and see discussion in Wasow, 2022) supports the opposing claim by Surprisal Theory that words with higher information content (or which improve predictability of following words) are more likely to occur first.
-Here the prediction is that a language with heavier nouns should display SV order, and one with heavier verbs should display VS order.
-Such predictions, if true, should be observable in crosslinguistic corpora.
-An important note here is that while the processing literature is largely grounded in a functional/usage-based perspective, it does not necessarily require a usage-based stance.
-That is, cognitive processes can be argued to apply at various levels of processing, including above the “innate” bedrock of current Minimalist theory (see Rastelli 2025).
-If there is a universal processing bias in the brain related to the complexity of structures or words in language, as has been claimed, such a bias would apply regardless of whether grammatical rules are innate or learned, and would impact or drive how such rules are realized or change over time.
-2.1 Main hypotheses
-Using word length as a measure of word complexity, the current paper compares average lengths of particular classes of words (nouns/arguments and verbs/predicates) with word order in individual languages.
-The two processing claims in the literature give rise to the following two competing hypotheses regarding this relationship:
+[^4] The procedure for determing the N1 ratio is essentially to count the number of sentences in a corpus that have both arguments (nouns, pronouns, etc) and predicates (verbs, auxiliaries), and then to derive a ratio from the number of argument-first sentences divided by the number of predicate-first sentences in that set.
+As an example, English [ISO 639-3: eng] has an N1 ratio of 4.51, whereas Irish [ISO 639-3: gle] has an N1 ratio of 1.24, based on their respective corpus in the taggedPBC.
+This ratio can then be related to other aspects of language, such as word order classifications.
+See Ring (2025) for further details.
+What this means is that using this dataset we can derive word order and word length information in a systematic manner, allowing for an investigation of word length in relation to word order.
+Figure 2: The N1 ratio and word order in 3 typological databases (p < 0.001; Ring 2025)
+The following studies focus on intransitive word order (i.e.
+the basic word order of intransitive clauses, in which a predicate requires a single “S” argument) in languages found in the taggedPBC.
+[^5] Minimal data and Python code to replicate the results in this paper can be found at https://osf.io/nsaq2 (3MB).
+More comprehensive data and code (7GB) can be found at the taggedPBC github repository (fn 2).
+The first study measures word classes, showing that average lengths of nouns and verbs in a language’s corpus are highly correlated with SV, VS, and free word order classifications when frequency is considered.
+The second study investigates whether this effect occurs outside of the taggedPBC, showing that information on average lengths of nouns and verbs computed from different corpora also allow for the classification of languages as SV or VS, differentiating between historical and modern languages for which word order changes are known.
+The third study examines whether the lengths of nouns and verbs in corpora provides a better indication of word order than known areal or genealogical signals, and shows that in heirarchical linear regressions, word length is indeed a better predictor of word order.
+2.1 Study 1: Word length and word order
+Study 1 follows previous studies in using word length as a measure of word complexity (Lewis & Frank 2016; Bentz et al.
+2016, 2017; Baerman et al.
+2015, 2017; Bentz et al.
+2023; and see Piantadosi et al., 2011; Meylan & Griffiths, 2021; Koplenig et al., 2022 for further discussion in relation to frequency in corpora), comparing average lengths of particular classes of words (nouns/arguments and verbs/predicates) with word order in individual languages.
+The two processing claims in the literature (from “efficiency” and “surprisal” approaches) give rise to the following two competing hypotheses regarding this relationship:
 H1:
@@
 (maximize complexity)
-2.2 Investigating word order and length
-Using the 1,597 languages in the tagged PBC, we can identify basic word order via the N1 ratio.
-[^7] As an example, English [ISO 639-3: eng] has an N1 ratio of 4.51 and is classed as SV, whereas Irish [ISO 639-3: gle] has an N1 ratio of 1.24 and is classed as VS.
-Details on the N1 ratio and its validity are given above and at the linked repository (Fn 4).
+For the 1,942 languages in the taggedPBC, we can identify basic word order via the N1 ratio.
 We can also identify the average length of nouns/arguments and verbs/predicates in each language’s corpus and see how well this correlates with the N1 ratio.
 2.1.1 Word order
-Of the languages in the tagged PBC, 976 are also found in typological databases with categorizations for word order.
-For the remainder, the work described at the linked repository has shown that the N1 ratio is a valid corpus-based proxy for word order.
-Accordingly I used the N1 ratio to impute the (SV, VS, free) category of the remaining 621 languages in the tagged PBC not categorized by WALS, AUTOTYP or Grambank.
-[^8] I trained a Gaussian Naive Bayes classifier on the known data to predict the category (SV, VS, free) of languages for which word order status was unknown, based on the N1 ratio as derived from the PBC.
-Code to reproduce these results can be found at the linked repository (Fn 4).
-This resulted in 1,251 languages classified as “SV”, 295 languages classified as “VS”, and 51 languages classified as “free”.
+Of the languages in the taggedPBC, 1,152 are also found in typological databases with categorizations for word order.
+For the remainder, the N1 ratio was used to impute the (SV, VS, free) category of the remaining 790 languages in the taggedPBC not categorized by WALS, AUTOTYP or Grambank, and the final categorizations were stored for all 1,942 languages.
+[^6] I trained a Gaussian Naive Bayes classifier on the known data to predict the category (SV, VS, free) of languages for which word order status was unknown, based on the N1 ratio as derived from the taggedPBC.
+Code to reproduce these results can be found at the linked repository (fn 2).
+This resulted in 1,548 languages classified as “SV”, 331 languages classified as “VS”, and 63 languages classified as “free”.
 2.1.2 Nouns and verbs
-To measure word length, I counted the number of characters in a word tagged as Noun/Argument or Verb/Predicate and computed an average length for each class on a per-language basis.
+To measure word length, I counted the number of characters in a word tagged as Noun/Argument or Verb/Predicate [^7] Here I conducted several tests to see if there were different effects for different word types and groupings.
+“Argument” is a grouping of words tagged as NOUN, PRON, PROPN in the universal tag set, while “predicate” refers to a grouping of VERB and AUX tags.
+and computed an average length for each class on a per-language basis.
 [^8] For example, English [ISO 639-3: eng] arguments have an average length of 5.9 characters, while English predicates have an average length of 5.75.
+English nouns have an average length of 5.9 characters, while English verbs have an average length of 5.8 characters.
 Irish [ISO 639-3: gle] arguments have an average length of 6.31 characters, while Irish predicates have an average length of 6.29 characters.
-In the tagged PBC corpus the average number of (unique) arguments per language is 732, while the average number of predicates is 495.
-The mean length of arguments is 6.76 characters, while the mean length of predicates is 6.90 characters.
-Conducting a Student’s T-Test showed that this is statistically significant (p = 0.006), indicating that crosslinguistically, arguments are shorter than predicates.
-All statistics showed a bimodal (normal) distribution.
+Irish nouns have an average length of 5.2 characters, while Irish verbs have an average length of 7.9 characters.
+In the taggedPBC the average number of (unique) arguments per language is 741, and the average number of predicates is 498.
+The mean length of arguments is 6.71 characters, and the mean length of predicates is 6.82 characters.
+Conducting a Student’s T-Test showed that this is statistically significant (p < 0.001), indicating that crosslinguistically, arguments are shorter than predicates.
+All statistics showed a normal distribution.
 2.1.3 Relationship between word order and noun/verb lengths
@@
 Analysis of the results of word order and its relation to length of the different argument and predicate categories is quite striking.
-First, both SV and VS languages have a highly significant effect of word length, [^10] Additional statistical comparisons and plots are available at the linked repository (Fn 4).
+First, both SV and VS languages have a highly significant effect of word length, [^9] Additional statistical comparisons and plots are available at the linked repository (Fn 2).
 whereby arguments are, overall, shorter than predicates crosslinguistically (Figure 3; p < 0.001).
@@
 This is a major finding that we will return to in the discussion.
-Setting it aside for now, we turn to two different tests of this measure relating to diachrony: predicting word order change, and the potential interaction with family membership (descent) and area (geography).
-2.3 Testing the predictive validity of word length
+Setting it aside for now, we turn to two different tests of this measure relating to diachrony: predicting word order change, and the potential interaction with family membership (genealogy/descent) and area (geography).
+2.2 Study 2: Predictive validity of word length
 Given the observation that relative lengths of nouns and verbs differ in corpora for languages with different word orders, we can ask how predictive this observation is.
@@
 However, we are also interested in understanding how noun and verb lengths may be related to word order.
-Being able to predict word order from noun and verb lengths in a corpus for both a modern language and its ancestral language would indicate that such properties are related to how word order evolves across time.
-To investigate this question, I sourced additional corpora of tagged data from two pairs of historical and modern languages in which the modern language changed its word order: 1) Ancient Hebrew (VS) and Modern Hebrew (SV) and 2) Classical Arabic (VS) and Egyptian Arabic (SV).
+Being able to predict word order only from noun and verb lengths in a corpus for both a modern language and its ancestral language would indicate that such properties are related to how word order evolves over time.
+To investigate this question, for Study 2 I sourced additional corpora of tagged data from two pairs of historical and modern languages in which the modern language changed its word order:
+  1. 1.
+    Ancient Hebrew (VS) and Modern Hebrew (SV)
+  2. 2.
+    Classical Arabic (VS) and Egyptian Arabic (SV)
 All languages belong to the Semitic branch of the Afroasiatic language family, a necessary constraint given the limited amount of data available for most pre-modern languages.
@@
 In the case of Classical Arabic I used the Quranic Arabic Corpus [^13] Version 0.4: https://corpus.quran.com and for Egyptian Arabic I used the BOLT Egyptian Arabic Treebank (Maamouri et al.
-2018).151515(Discussion Forum) https://catalog.ldc.upenn.edu/LDC2018T23 If the lengths of nouns and length of verbs in these corpora allows us to correctly classify Ancient Hebrew and Classical Arabic as VS and the modern varieties as SV, this would suggest that these are salient properties affecting how word order has developed diachronically.
+2018).
+[^14] https://catalog.ldc.upenn.edu/LDC2018T23 (Discussion Forum) If the lengths of nouns and verbs in these corpora allows us to correctly classify Ancient Hebrew and Classical Arabic as VS and the modern varieties as SV, this would suggest that these are salient properties affecting how word order has developed diachronically.
 Importantly, given the textual differences between the taggedPBC (New Testament texts) and the test corpora (Biblical, Quranic, modern, and informal texts), such a finding would be significant.
@@
 Accordingly, I use the abbreviation cla here, as there is no current language assigned to this code in the ISO 639-3 tables.
-and Egyptian Arabic (ISO 639-3: arz) were processed and the respective word classes were counted and measured as per the method outlined for the tagged PBC (each corpus > 400 sentences, see Table 1).
+and Egyptian Arabic (ISO 639-3: arz) were processed and the respective word classes were counted and measured as per the method outlined for Study 1 (each corpus > 400 sentences, see Table 1).
 The average length of nouns and verbs (including frequency effects) were then used as continuous features for a Gaussian Naive Bayes (GNB) classifier.
@@
 The third measure is the ratio of noun and verb lengths, and is related to maximization of complexity (via “entropy”), whereby a lower ratio indicates that the particular word classes are less likely to be easily differentiable (and therefore predictable) based on their length.
-Using these five features, a GNB classifier was trained on the original combined word order dataset containing hand-coded classification of 974 languages (Modern Hebrew and Modern Arabic were removed), and then used to predict word order for our pairs of languages: Ancient Hebrew and Modern Hebrew on the one hand, Classical Arabic and Egyptian Arabic on the other.
+Using these five features (nounlen, verblen, nlonger, nlen-vlen, nlen/vlen), a GNB classifier was trained on the original combined word order dataset from the taggedPBC containing hand-coded classification of 1,150 languages (Modern Hebrew and Modern Arabic were removed), and then used to predict word order for our pairs of languages: Ancient Hebrew and Modern Hebrew on the one hand, Classical Arabic and Egyptian Arabic on the other.
 Ancient Hebrew was classified as VS while Modern Hebrew was classified as SV by the trained classifier.
 Classical Arabic was classified as VS while Egyptian Arabic was classified as SV.
-This indicates that the relative lengths of nouns and verbs, when frequency is taken into account, can predict word order from corpora, with various implications for our understanding of language processing and language change.
+As an additional check, I sourced data from another historical language, Old Irish, as represented in the Codex Paulinus Wirziburgensis, whereby Latin text of St.
+Paul’s epistles was annotated by Irish scholars during the middle of the eighth century.
+The dataset (Table 2; Doyle 2018) contains over 3,600 Irish glosses/sentences.
+As an Indo-European language, Old Irish is widely surmised to have changed its word order to verb-initial from an earlier state, so this check allows us to verify whether the same properties hold in relation to lengths of nouns and verbs.
+Indeed, the classifier accurately classified Old Irish as VS, based on the data extracted from the Würzburg Glosses.
+Language ISO 639-3 Data Source Sentences Unique nouns Unique verbs
+Old Irish sga Würzburg 3,647 743 435
+Table 2: Irish corpus used for testing historical word order change
+Study 2 thus indicates that the relative lengths of nouns and verbs, when frequency is taken into account, can predict word order from corpora unrelated to the type of corpora contained in the taggedPBC, with various implications for our understanding of language processing and language change.
 This is a striking finding, and I am unaware of any other research that has been able to show predictive validity of a model related to word order change.
-However, it is possible that this result is conditioned to some degree by the fact that these are Semitic languages spoken in the same geographical area, so we now turn to an investigation of whether descent or location provides a better explanation for the correlation of noun and verb lengths with word order.
-2.4 Testing variance predicted by word length
+However, it is possible that this result is conditioned to some degree by the fact that the majority of the test cases are Semitic languages spoken in the same geographical area, so we now turn to an investigation of whether genealogical descent or location provides a better explanation for the correlation of noun and verb lengths with word order.
+2.3 Study 3: Testing variance predicted by word length
 Dunn et al.
 (2011) showed that descent from a common ancestor was a better predictor of many grammatical patterns than other explanations.
-Although Jäger & Wahle (2021) counter this finding with additional data, we can also assess this claim using the current dataset to see if descent is more explanatory of word order for a given language than the lengths of nouns and verbs.
-Additionally, in more recent work, Naranjo & Becker (2022) show that areal affects are as important as family membership in predicting linguistic phenomena, finding that the results in Dunn et al can be just as easily predicted by language area.
+Although Jäger & Wahle (2021) counter this claim with additional data, we can also use the current dataset to see if descent is more explanatory of word order for a given language than the lengths of nouns and verbs.
+Additionally, in more recent work, Guzmán Naranjo & Becker (2022) show that areal affects are as important as family membership in predicting linguistic phenomena, finding that the results in Dunn et al can be just as easily predicted by language area.
 This is also a claim that we can test with our dataset.
-Accordingly, this investigation can proceed with several (motivated) simplifying observations.
+Accordingly, Study 3 can proceed with several (motivated) simplifying observations.
 First, for each language we can take the highest level node of family membership that is known based on current scholarship (via Glottolog).
@@
 Finally, we can remove languages with “free” word order from consideration, as this leaves us with a binary distinction in word order (SV, VS) that opens up more possibilities for statistical tools.
-In particular, a binary dependent variable for word order means we can use hierarchical linear regression, which allows us to observe the relative impact of particular features that are theoretically relevant to the question at hand by seeing how they influence the performance of successive models.
+In particular, a binary dependent variable for word order means we can use hierarchical linear regression (HLR), which allows us to observe the relative impact of particular features that are theoretically relevant to the question at hand by seeing how they influence the performance of successive models.
 In the case of word order, we can train a regression model with a feature known to be relevant to word order (the N1 ratio), and then subsequent models that incorporate additional features of interest (“language area/location”, “family membership”, “noun/verb lengths”).
@@
 Accordingly, for all languages in the taggedPBC I retrieved the family membership and geographical information from Glottolog, filling in missing data via Ethnologue and other sources.
-Then I filtered the data on four conditions: a) languages found in Dunn et al that are also present in the tagged PBC (106 languages), b) languages found in the tagged PBC that are in the four families investigated by Dunn et al (539 languages), c) languages in families with more than 75 members in the tagged PBC (962 languages), and d) languages in families with 2 or more members in the tagged PBC (1,486 languages).
-For each of these conditions, I ran a hierarchical linear regression, adding features of interest progressively to see how they impacted the fit of the successive models and whether particular features account for additional variance.
+Then I filtered the data on four conditions: a) languages found in Dunn et al that are also present in the taggedPBC (120 languages), b) languages found in the taggedPBC that are in the four families investigated by Dunn et al (662 languages), c) languages in families with more than 75 members in the taggedPBC (1182 languages), and d) languages in families with 2 or more members in the taggedPBC (1,805 languages).
+For each of these conditions, I ran an HLR, adding features of interest progressively to see how they impacted the fit of the successive models and whether particular features account for additional variance.
 The base model (Model 1) uses the N1 ratio as the primary feature for classification.
-This is because we know from comparison with expert determinations of word order (see footnote 5 above) that the N1 ratio significantly correlates with the classification of languages as SV/VS.
-Due to the strong relationship observed between language area and typological properties of language (see Naranjo & Becker 2022), Model 2 then includes GPS coordinates and macroarea identifiers for each language as features.
+This is because we know from comparison with expert determinations of word order (see Ring 2025) that the N1 ratio significantly correlates with the classification of languages as SV or VS.
+Due to the strong relationship observed between language area and typological properties of language (see Guzmán Naranjo & Becker 2022), Model 2 then includes GPS coordinates and macroarea identifiers for each language as features.
 Model 3 additionally incorporates descent (language family membership) as a feature (as per Dunn et al.
@@
 a) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
-1 (N1 ratio) 106 45.85 7.8e-10
-2 (Lg area) 106 20.40 2.4e-12 8.58 3.9e-05**
-3 (Family) 106 16.70 5.8e-12 1.48 0.227
-4 (N/V len) 106 13.40 4.9e-12 3.27 0.042*
+1 (N1 ratio) 120 45.27 6.4e-10
+2 (Lg area) 120 24.29 1.3e-14 12.78 2.87e-07 **
+3 (Family) 120 20.02 2.7e-14 2.05 0.154
+4 (N/V len) 120 15.75 2.9e-14 3.17 0.045 *
 b) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
-1 (N1 ratio) 539 230.41 1.4e-43
-2 (Lg area) 539 100.00 1.7e-63 39.86 2.9e-23**
-3 (Family) 539 80.71 5.3e-63 2.46 0.118
-4 (N/V len) 539 64.75 3.5e-67 14.57 6.9e-07**
+1 (N1 ratio) 662 189.99 3.6e-38
+2 (Lg area) 662 95.65 4.1e-64 50.07 3.7e-29 **
+3 (Family) 662 78.23 2.4e-64 5.79 0.016 *
+4 (N/V len) 662 65.43 2.6e-71 21.30 1.0e-09 **
 c) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
-1 (N1 ratio) 962 426.74 9.9e-79
-2 (Lg area) 962 220.49 4.2e-134 105.35 6.1e-59**
-3 (Family) 962 176.22 6.6e-133 0.04 0.839
-4 (N/V len) 962 130.93 8.3e-135 9.69 6.7e-05**
+1 (N1 ratio) 1182 378.83 2.1e-73
+2 (Lg area) 1182 210.38 3.3e-136 116.99 2.6e-66 **
+3 (Family) 1182 169.09 1.3e-135 2.72 0.099
+4 (N/V len) 1182 127.47 2.3e-139 14.03 9.4e-07 **
 d) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
-1 (N1 ratio) 1486 586.46 1.8e-109
-2 (Lg area) 1486 193.61 1.3e-133 45.19 6.0e-28**
-3 (Family) 1486 155.10 1.2e-132 1.03 0.309
-4 (N/V len) 1486 115.43 3.5e-135 11.02 1.7e-05**
+1 (N1 ratio) 1805 535.41 6.1e-104
+2 (Lg area) 1805 178.89 3.8e-129 46.52 5.8e-29 **
+3 (Family) 1805 143.49 2.3e-128 1.66 0.197
+4 (N/V len) 1805 108.16 2.2e-132 14.46 5.8e-07 **
 Table 3: Hierarchical Linear Regressions under 4 conditions
@@
 Additionally, language area is a significant predictor of word order that adds to the strength of the classification.
-Language family (descent) does not add additional predictive power to our model under any conditions (over language area), whereas the length of nouns and verbs does.
+Language family (descent) does not add additional predictive power to our model under most conditions (over language area), whereas the length of nouns and verbs does.
 This analysis shows that the length of nouns and verbs in a language accounts for more variance and at the same time is more significant than that language’s family membership in determining its word order classification, under all test conditions.
-When we remove language area from the linear regression this is still the case, [^18] See additional statistical results at the linked repository (Fn 4).
+When we remove language area from the linear regression this is still the case, [^17] See additional statistical results at the linked repositories (Fn 2; Fn 5).
 indicating that while language area is a stronger predictor than family membership, lengths of nouns and verbs is yet a stronger one.
@@
 If a large number of languages in a sample are not related to others, clearly descent can have only a minimal impact on word order.
-By including only language families with two or more members (which excludes 111 languages in the taggedPBC), I show that descent does not explain much variance above and beyond language location/area, in contrast with noun and verb lengths.
+By including only language families with two or more members in the taggedPBC (which excludes 111 languages), I show that descent does not explain much variance above and beyond language location/area, in contrast with noun and verb lengths.
 This also suggests that language family/descent is highly correlated with linguistic area, which conforms with expectations related to human history and migration.
@@
 3 Discussion
-That arguments overall are shorter in languages of the tagged PBC, across all word orders, suggests that there are certain (non-Noun) arguments affecting our observations.
+That arguments overall are shorter in languages of the taggedPBC, across all word orders, suggests that there are certain (non-Noun) arguments affecting these observations.
 Particularly, the effect of PRON seems quite likely, whereby a tendency for pronouns to be shorter crosslinguistically, along with a preference for pronouns as arguments, may impact the observation that arguments are shorter overall than predicates.
-This pattern aligns with the processing claim that ‘heavier’ words (or word classes in this case) occur later in a sentence, partly validating Hypothesis 1.
-The counter-argument, also from the data, is that nouns are longer than verbs in SV languages and shorter than verbs in VS languages when frequency is considered, validating Hypothesis 2.
+This pattern aligns with the processing claim that “heavier” words (or word classes in this case) occur later in a sentence, partly validating Hypothesis 1 of study 1.
+The counter-argument, also from the data, is that nouns are longer than verbs in SV languages and shorter than verbs in VS languages when frequency is considered, validating Hypothesis 2 of study 1.
 This suggests that speakers prefer to place heavier (more “contentful”, or perhaps “novel”) constituents before lighter ones crosslinguistically in many cases.
-These seemingly antithetical results partially validate both the “weight”-based theory and the “surprisal” theory, highlighting the interaction of two pressures: one whereby speakers prioritize lighter constituents (possibly for the purposes of production/processing efficiency), and one where they prioritize heavier constituents (possibly for informational purposes).
+These seemingly antithetical results partially validate both the “efficiency”-based theory and the “surprisal” theory, highlighting the interaction of two pressures: one whereby speakers prioritize lighter constituents (possibly for the purposes of production/processing efficiency), and one where they prioritize heavier constituents (possibly for informational purposes).
 The two perspectives may in fact be looking at the same data from different angles.
@@
 The current proposal re-formulates or refines Hawkins’ PGCH to incorporate (increased) complexity as a motivation for the order of constituents, bearing some resemblance to proposals in Levshina (2023).
-I frame this as (to borrow a term from computer gaming) the “Min-Max” theory of processing (schematized in Figure 3).
-The idea here is that speakers seek to minimize their processing effort while maximizing the information in their utterances.
+I describe this as (to borrow a term from computer gaming) [^19] A “min-maxer” in gaming terminology is an individual who minimizes their effort while maximizing their gains.
+This can be in terms of focusing their efforts on activities in a game that give the best returns, or by minimizing survival stats on a character in order to maximize damage (as in a “glass cannon” build for ranged attackers in an rpg).
+the “Min-Max” theory of processing (schematized in Figure 5), whereby speakers seek to minimize their processing effort while maximizing the information in their utterances.
 These are competing/parallel pressures that speakers are constantly balancing in production (or “performance”) in order to choose the best strategy for their immediate goals.
@@
 In a study of patients with nonfluent aphasia, Rezaii et al.
-(2023) showed that patients with this condition compensate for their deficit in producing complex grammatical structures by increasing the information content of their utterances, dropping high-frequency (function) words in favor of low-frequency (content) words.
-By modeling variations in grammatical structure based on corpora from 80 languages, Hahn & Xu (2022) found that the resulting models were less optimized for dependency length and information locality than the original structures of each language, indicating that languages evolve to efficiently balance the two considerations.
+(2023) showed that patients with this condition compensate for their deficit in producing complex grammatical structures by increasing the information content of their utterances, dropping high-frequency (function) words in favor of low-frequency (content) words (i.e.
+“maximizing information”).
+By modeling variations in grammatical structure based on corpora from 80 languages, Hahn & Xu (2022) found that the resulting models were less optimized for dependency length (DL) and information locality (IL) than the original structures of each language, suggesting that languages evolve to efficiently balance the two considerations of DL and IL.
 [^20] Notably, Hahn & Xu propose a theory that the interaction between Dependency Length (DL) and Information Locality (IL) accounts for word order variation and change, but neither of the two concepts, as expressed, make claims about directionality/order, only about the relationships between (internal) constituents.
 In contrast, the Min-Max theory makes the claim that speakers balance information content (as reflected by longer nouns in SV languages and longer verbs in VS languages) with processing/production concerns, and that this interaction is what drives the evolution of word order.
-As nouns lengthen or verbs shorten in a language, this will affect their placement in a stream, which then impacts the overall order observed in surface structures.
+As nouns lengthen or verbs shorten in a language, this will affect their placement in a stream, which then impacts the overall linear order observed in surface structures.
 That is, the Min-Max theory makes a universal claim regarding language-external pressures, while DL and IL are language-internal concepts.
-These and other studies indicate that language users are sensitive to efficiency as it pertains both to processing and information content.
+These and other studies indicate that language users are sensitive to the pressures of minimizing processing and maximizing information content.
 The sensitivity of language users to processing efficiency and information content has also been a major topic of investigation in information theory, which is concerned in part with the encoding and decoding of information streams.
@@
 minimizing the representational complexity, and maximizing the relevant information about Y” (Zaslavsky 2020: 14).
-Work by Zaslavsky shows that IB can account for the development of color naming systems in language among other kinds of sytems, and additional research shows the utility of IB in machine learning and distributional models of semantics (Slonim & Tishby 2000).
-The implications of this principle have most recently been examined in agent-based emergent learning simulations with the addition of a “utility” principle (Tucker et al.
+Zaslavsky shows that IB can account for the development of color naming systems in language, and additional research shows the utility of IB in machine learning and distributional models of semantics (Slonim & Tishby 2000).
+Implications of this interaction have most recently been examined in agent-based emergent learning simulations with the addition of a “utility” principle (Tucker et al.
 2025) – the interaction between “utility”, “informativeness”, and “complexity” allows more human-like communication to emerge among agent models in their study.
@@
 Speakers min-max in order to select words in a stream (“informativeness” and “complexity” as per Tucker et al.), while simultaneously min-maxing to select relevant structures related to their communicative goals (Tucker et al.’s “utility”).
-The Min-Max theory also bears some resemblance to Levshina’s (2023) proposals, described in terms of communicative efficiency and accessibility.
+As noted above, the Min-Max theory also bears some resemblance to Levshina’s (2023) proposals, described in terms of communicative efficiency and accessibility.
 However, such indicators cannot fully explain certain patterns in grammar such as “harmonic” word order.
-A recent proposed explanation for such patterns involves frequency of word classes, whereby Mansfield & Krapp (2025) suggest that “harmonic ordering is really a frequency effect: as replication converges on a phrase structure with a consistent linear order, the most frequent word class tends to be at one edge.” Although they did not test token length in their model, it is notable that in the current study frequency is a major factor in identifying the relationship between length and word order, suggesting that the most frequent word class in their model may also be the longest element of the domain they are testing, which would be driven by a maximization of information principle.
-The Min-Max theory also seems to provide an explanation for non-parallel (hierarchical) harmonic word order, which is not explained by their model.
-The Min-Max theory may additionally provide an explanation for why word forms remain long (or lengthen), as described for particular languages (see Cristofaro 2021; Haspelmath 2021; Petré 2017).
+A recent proposed explanation for such patterns involves frequency of word classes, whereby Mansfield & Krapp (2025) suggest that “harmonic ordering is really a frequency effect: as replication converges on a phrase structure with a consistent linear order, the most frequent word class tends to be at one edge.” Although they did not test token length in their model, it is notable that in Study 1 (§2.1 above) frequency is a major factor in identifying the relationship between length and word order, suggesting that the most frequent word class in their model may also be the longest element of the domain they are testing, which would be driven by the maximization of information principle.
+The Min-Max theory might also provide an explanation for non-parallel (hierarchical) harmonic word order, which is not explained by their model.
+The Min-Max theory may additionally explain why word forms remain long (or lengthen), as described for particular languages (see Cristofaro 2021; Haspelmath 2021; Petré 2017).
 Here, again, the maximization of information principle would support the development or retention of particular (long) word forms, as speakers balance this need against the minimization of processing difficulty.
-As noted above, this proposal is supported by evidence from color-naming systems, which show patterns in keeping with the Information Bottleneck principle.
+As noted above, this proposal is supported by evidence from color-naming systems, which show patterns in keeping with IB (Zaslavsky 2020).
 The Min-Max theory would suggest that this principle does not only hold for “forms” (of, i.e.
-color words), but also governs “behavior”: speakers behave in ways that are consistent with maximizing information in streams (via word order), while minimizing processing difficulty.
-As noted in section 2, such a theory can be conceived of as a universal (underlying) pressure that applies generally to how language is realized.
-From an innate perspective, speaker choices impacted by such pressures would lead to a re-analysis or re-formation of rules that guide surface realizations.
+color words), but also governs “behavior”: speakers behave in ways that are consistent with maximizing information in streams (via word order), while minimizing processing effort.
+Such a theory can be conceived of as a universal (underlying) pressure that applies generally to how language is realized.
+From an innateness perspective, speaker choices impacted by such pressures would lead to a re-analysis or re-formation of rules that guide surface realizations.
 From a usage-based perspective such pressures would lead to the development of new surface variants.
@@
 This indicates that (maximizing) information content is more important to speakers than (minimizing) processing complexity.
-This could explain the “agent/subject-first preference” in experiments (Sauppe et al.
-2023; Culbertson et al.
-2012), as the “information maximization” pressure wins out in cases where this most efficiently meets communicative goals.
+This could explain the “agent/subject-first preference” in experiments (Culbertson et al.
+2012; Sauppe et al.
+2023), as the “information maximization” pressure wins out in cases where this most efficiently meets communicative goals.
 Further, the data shows that all languages have a relatively high proportion of sentences in which the argument occurs before the predicate.
@@
 Following on from Zipf’s and subsequent observations, there is no doubt that the length of strings is an important factor in processing streams, but clearly not all strings are created or treated equally by speakers.
-This study shows that nouns are treated differently (or at least have a different distributional relationship) than other kinds of arguments, which leads to different realizations of word order crosslinguistically.
-The pressures identified here have wide-ranging implications and applications for any conceptualization of how language operates, whether in terms of understanding the intricacies of individual linguistic scenarios, designing intelligent systems that use language, or developing clinical interventions to assist speakers in various ways.
-3.2 Limitations
-One limitation in this study is the use of automated processes to identify part of speech tags in the data.
+The studies presented here show that nouns are treated differently (or at least have a different distributional relationship) than other kinds of arguments, which leads to different realizations of word order crosslinguistically.
+The pressures identified in this paper have wide-ranging implications and applications for any conceptualization of how language operates, whether in terms of understanding the intricacies of individual linguistic scenarios, designing intelligent systems that use language, or developing clinical interventions to assist speakers in various ways.
+3.2 Limitations and future directions
+One limitation in this set of studies is the use of automated processes to identify part of speech tags in the data.
 Ideally, all the data should be hand-annotated for parts of speech, which would allow for better comparison of the different word categories across languages.
@@
 (2023).
-The issue of word complexity also could use further expansion to cover a broader range of possibilities than simply the length of a word based on a set of (roman) characters.
+The issue of word complexity could additionally use further expansion to cover a broader range of possibilities than simply the length of a word based on a set of (roman) characters.
 A final note is that our conclusions are only as good as the data that we have available to us.
@@
 On the basis of this investigation I have proposed a parsimonious “Min-Max” theory of processing and information structure that underlies language evolution, whereby speakers minimize processing effort while maximizing information gain.
-This proposal is in keeping with at least one existing information-theoretic proposal (IB) as well as theories of processing (PGCH; surprisal) and can explain the correlation between the length of nouns/verbs and basic word order, given that these features allow us to predict the word order of a language.
+This proposal is in keeping with at least one existing information-theoretic proposal (IB) as well as theories of processing (efficiency; PGCH; surprisal) and can explain the correlation between the length of nouns/verbs and basic word order, given that these features allow us to predict the word order of a language.
 It is as yet unclear whether such a principle would apply at multiple hierarchical levels within language, which is an area that deserves further careful investigation with datasets such as the taggedPBC.

### v3

@@
 2011; Wasow 2022; Levy 2008) that make different predictions about how word order should be realized crosslinguistically.
-The marriage of both “efficiency” and “surprisal” approaches under the Min-Max theory is justified with evidence from a massive dataset of 1,942 language corpora tagged for parts of speech (Ring 2025), in which average lengths of particular word classes correlates with word order, allowing for prediction of basic word order from diverse corpora.
-The general universal pressure of word class length in corpora is shown to give a stronger explanation for word order realization than either genealogical or areal factors, highlighting the importance of language corpora for investigating such questions.
+The marriage of both “efficiency” and “surprisal” approaches under the Min-Max theory is justified with evidence from a massive dataset of 1,942 language corpora tagged for parts of speech (Ring 2026), in which average lengths of particular word classes correlates with word order, allowing for prediction of basic intransitive word order from diverse corpora.
+The general universal pressure of word class length in corpora is shown to give a stronger explanation for word order realization than either genealogical or areal factors, highlighting the importance of language corpora for investigating such questions
 1 Introduction
@@
 The first (typological) approach is to observe word orders in natural languages and identify patterns whereby the orders of certain constituents are correlated with other grammatical patterns.
-This approach has developed multiple crosslinguistic databases with expert determinations of word order (WALS, Dryer & Haspelmath 2013; AUTOTYP, Bickel et al.
+This approach has developed multiple crosslinguistic databases with expert determinations of word order (such as WALS, Dryer & Haspelmath 2013; AUTOTYP, Bickel et al.
 2023; Grambank, Skirgård et al.
@@
 2023), “efficiency” (Zipf, 1949; Wasow, 2002; Futrell et al., 2015b), and “surprisal” (Hale 2001; Levy 2008).
-While the typological observations do not make predictions about a language’s basic word order (the universals take basic word orders as a starting point and make further predictions, see Verkerk et al.
+While the typological observations do not make predictions about a language’s basic word order (the statistical universals take basic word orders as a starting point and make further predictions, see Verkerk et al.
 2025), the processing research does imply some particular claims and predictions about how words should be ordered in languages of the world.
@@
 2022).
-At the same time, findings from language processing studies suggest that other pressures mediate the realization of word order patterns, namely in relation to “efficiency” and “surprisal”.
+At the same time, findings from additional language processing research suggests that other pressures mediate the realization of word order patterns, namely in relation to “efficiency” and “surprisal”.
 The efficiency literature largely follows from Zipf’s (1932; 1949) observations that a) languages show a logarithmic distribution whereby the most frequently used words occur much more often than the next most frequent, and that b) the most frequent words in a language are also the shortest.
-The observation that shorter (or less complex) elements are more likely to occur initially in a sentence, while longer (or more complex) elements occur later, can be traced to Behaghel (1909), and the general idea is that this supports efficient communication.
+The finding that shorter (or less complex) elements are more likely to occur initially in a sentence, while longer (or more complex) elements occur later, is one that can be traced to Behaghel (1909), and the general idea is that this supports efficient communication.
 In contrast to this view, the surprisal literature focuses on how salient or predictable items are, suggesting that longer (or more complex) words are more likely to occur first in a sentence since they improve predictability of following constituents (Hale 2001; Levy 2008; Sauppe et al.
@@
 The lack of testing highlights the complexity of investigating pressures that influence the evolution and realization of language structures, as well as the dearth of datasets and databases that could address these questions.
-[^1] Datasets that have been used by previous research include the Universal Dependencies Treebanks (Zeman et al.
+[^1] Datasets that have been used by previous research include the Universal Dependencies Treebanks (UDT; Zeman et al.
 2024) and ODIN (Xia et al.
@@
 In particular, questions regarding the length of constituents in relation to their order in sentences requires a large number of sentences tagged for parts of speech (at minimum).
-Given that languages develop over time via cultural evolution, it is also difficult to separate general cognitive pressures from historical (genealogical) or areal factors.
+Given that languages develop over time via cultural evolution, it is also difficult to separate general cognitive pressures from historical (genealogical or areal) factors.
 This means that inferring crosslinguistic or universal principles necessitates both a large number of languages as well as sufficient diversity of languages to establish statistical validity.
@@
 The current paper takes advantage of a massive tagged parallel corpus to show how these pressures are realized in actual language data in relation to word order.
-While the dataset is still under development, the baseline data is tagged for parts of speech, which allows us to derive a proxy measure for word order from individual language corpora, as well as to measure other linguistic properties such as the average length of various word types.
+While the dataset is still under development, the baseline data is tagged for parts of speech, which allows us to derive a proxy measure for word order from individual language corpora.
+It also allows for measurement of other linguistic properties such as the average length of various word types, while also controlling for potential genealogical and areal bias.
 Results show that basic word order in intransitive clauses is highly correlated with the length of particular word classes, allowing for the prediction of a language’s word order based on corpora not in the dataset and explaining variance better than both genealogy and areal factors.
 In the remainder of this paper I first highlight the properties of the dataset (§2).
-I then discuss three studies that investigate word length in relation to word order (§2.1), the predictive validity of word lengths for word order (§2.2), and the variance predicted by word length in regression models (§2.3).
-I conclude with a discussion of the implications of these studies in relation to language evolution (§3), presenting the Min-Max theory of language behavior (§3.1) as well as some limitations and future directions regarding the current effort (§3.2).
+I then present three studies that investigate word length in relation to word order (§2.1), the predictive validity of word lengths for word order (§2.2), and the variance predicted by word length in regression models over and above language family and linguistic area (§2.3).
+I conclude with the implications of these studies in relation to language evolution (§3), discussing how the Min-Max theory of language behavior unites competing pressures (§3.1), as well as some limitations and future directions regarding the current effort (§3.2).
 2 Methodology and hypotheses
-The taggedPBC (Ring 2025) [^2] https://github.com/lingdoc/taggedPBC - see fn 5 for a link to a minimal OSF repository that allows for replication of the studies in this paper.
+The taggedPBC (Ring 2026) [^2] https://github.com/lingdoc/taggedPBC - see fn 5 for a link to a minimal OSF repository that allows for replication of the studies in this paper.
 contains pos-tagged parallel text data in 1,942 languages (roughly 25% of the world’s languages), with over 1,800 parallel verses for more than 1,850 of those languages.
 The dataset represents 155 language families and 78 isolates, dwarfing previously available resources, and the tagging methodology (via crosslingual tag transfer) has been shown to correspond well with existing automated taggers for high-resource languages (as found in the SpaCy and TranKit NLP libraries).
-[^3] https://spacy.io ; https://nlp.uoregon.edu/trankit Additionally, a novel measure extracted from individual corpora in this dataset (the “N1 ratio”) is significantly correlated (p < 0.001) with expert determinations of a language’s basic word order for intransitive clauses (“SV”, “VS”, “free”; Figure 2) as identified in the typological databases WALS (Dryer & Haspelmath 2013), Grambank (Skirgård et al.
-2023), and Autotyp (Bickel et al.
-2023).
-[^4] The procedure for determing the N1 ratio is essentially to count the number of sentences in a corpus that have both arguments (nouns, pronouns, etc) and predicates (verbs, auxiliaries), and then to derive a ratio from the number of argument-first sentences divided by the number of predicate-first sentences in that set.
+[^3] https://spacy.io ; https://nlp.uoregon.edu/trankit Additionally, a novel measure extracted from individual corpora in this dataset (the “N1 ratio”) significantly correlates (p < 0.001) with expert determinations of a language’s basic word order for intransitive clauses (“SV”, “VS”, “free”; Figure 2) as identified in the typological databases WALS, Grambank, and AUTOTYP.
+[^4] The procedure for determing the N1 ratio is essentially to count the number of sentences in a corpus that have both arguments (nouns, pronouns, proper nouns) and predicates (verbs, auxiliaries), and then to derive a ratio from the number of argument-first sentences divided by the number of predicate-first sentences in that set.
 As an example, English [ISO 639-3: eng] has an N1 ratio of 4.51, whereas Irish [ISO 639-3: gle] has an N1 ratio of 1.24, based on their respective corpus in the taggedPBC.
@@
 What this means is that using this dataset we can derive word order and word length information in a systematic manner, allowing for an investigation of word length in relation to word order.
-Figure 2: The N1 ratio and word order in 3 typological databases (p < 0.001; Ring 2025)
+Figure 2: The N1 ratio and word order in 3 typological databases (p < 0.001; Ring 2026)
 The following studies focus on intransitive word order (i.e.
@@
 More comprehensive data and code (7GB) can be found at the taggedPBC github repository (fn 2).
-The first study measures word classes, showing that average lengths of nouns and verbs in a language’s corpus are highly correlated with SV, VS, and free word order classifications when frequency is considered.
-The second study investigates whether this effect occurs outside of the taggedPBC, showing that information on average lengths of nouns and verbs computed from different corpora also allow for the classification of languages as SV or VS, differentiating between historical and modern languages for which word order changes are known.
-The third study examines whether the lengths of nouns and verbs in corpora provides a better indication of word order than known areal or genealogical signals, and shows that in heirarchical linear regressions, word length is indeed a better predictor of word order.
+The first study (§2.1) measures word classes, showing that average lengths of nouns and verbs in a language’s corpus are highly correlated with SV, VS, and free word order classifications when frequency is considered, and that this holds even when controlling for family and areal bias.
+The second study (§2.2) investigates whether this effect occurs outside of the taggedPBC, showing that information on average lengths of nouns and verbs computed from different corpora also allow for the classification of languages as SV or VS, differentiating between historical and modern languages for which word order changes are known.
+The third study (§2.3) examines whether the lengths of nouns and verbs in corpora provides a better indication of word order than known areal or genealogical signals, and shows that in heirarchical linear regressions, word length is indeed a better predictor of word order.
 2.1 Study 1: Word length and word order
-Study 1 follows previous studies in using word length as a measure of word complexity (Lewis & Frank 2016; Bentz et al.
+Study 1 follows previous research in using word length as a measure of word complexity (Lewis & Frank 2016; Bentz et al.
 2016, 2017; Baerman et al.
@@
 2023; and see Piantadosi et al., 2011; Meylan & Griffiths, 2021; Koplenig et al., 2022 for further discussion in relation to frequency in corpora), comparing average lengths of particular classes of words (nouns/arguments and verbs/predicates) with word order in individual languages.
+[^6] In the following studies romanized orthographic characters are used as a proxy for phonemes.
 The two processing claims in the literature (from “efficiency” and “surprisal” approaches) give rise to the following two competing hypotheses regarding this relationship:
@@
 The average length of nouns is shorter in languages classed as Subject-initial, while the average length of verbs is shorter in languages classed as Verb-initial.
-(minimize complexity)
+(efficiency - minimize complexity)
 H2:
@@
 The average length of nouns is longer in languages classed as Subject-initial, while the average length of verbs is longer in languages classed as Verb-initial.
-(maximize complexity)
+(surprisal - maximize complexity)
 For the 1,942 languages in the taggedPBC, we can identify basic word order via the N1 ratio.
@@
 Of the languages in the taggedPBC, 1,152 are also found in typological databases with categorizations for word order.
-For the remainder, the N1 ratio was used to impute the (SV, VS, free) category of the remaining 790 languages in the taggedPBC not categorized by WALS, AUTOTYP or Grambank, and the final categorizations were stored for all 1,942 languages.
+For the remainder, the N1 ratio was used to impute the (SV, VS, free) category of the remaining 790 languages in the taggedPBC not categorized by WALS, Grambank or AUTOTYP, and the final categorizations were stored for all 1,942 languages.
 [^7] I trained a Gaussian Naive Bayes classifier on the known data to predict the category (SV, VS, free) of languages for which word order status was unknown, based on the N1 ratio as derived from the taggedPBC.
@@
 This resulted in 1,548 languages classified as “SV”, 331 languages classified as “VS”, and 63 languages classified as “free”.
-2.1.2 Nouns and verbs
-To measure word length, I counted the number of characters in a word tagged as Noun/Argument or Verb/Predicate [^7] Here I conducted several tests to see if there were different effects for different word types and groupings.
-“Argument” is a grouping of words tagged as NOUN, PRON, PROPN in the universal tag set, while “predicate” refers to a grouping of VERB and AUX tags.
+2.1.2 Measuring word length by class
+To measure word length, I counted the number of characters in a word tagged as Noun/Argument or Verb/Predicate [^8] Here I conducted several tests to see if there were different effects for different word types and groupings, where “argument” is a grouping of words tagged as NOUN, PRON, PROPN in the universal tag set, while “predicate” refers to a grouping of VERB and AUX tags.
 and computed an average length for each class on a per-language basis.
@@
 All statistics showed a normal distribution.
+Since we are working with corpus data, there are two different length measures that we can compute for word classes in each language.
+We can compute average lengths for classes based on 1) the unique items in that class, and 2) the total items in the class.
+The first measure is similar to the information that can be gleaned from a wordlist or dictionary, where infrequent terms have the same weight as frequent items.
+The second measure gives more weight to frequent items.
+Since we know that frequency of occurrence in corpora is a factor in constituent predictability (Piantadosi et al., 2011; Meylan & Griffiths, 2021; Koplenig et al., 2022), it is worth observing whether the two measures differ in terms of their relationship with word order.
 2.1.3 Relationship between word order and noun/verb lengths
-To see whether word order is related to the length of nouns and verbs I conducted a repeated-measures ANOVA with the average lengths of nouns and verbs as the repeated measures and the word order (SV, VS, free) as the between subject factors.
-Analysis of the results of word order and its relation to length of the different argument and predicate categories is quite striking.
-First, both SV and VS languages have a highly significant effect of word length, [^9] Additional statistical comparisons and plots are available at the linked repository (Fn 2).
-whereby arguments are, overall, shorter than predicates crosslinguistically (Figure 3; p < 0.001).
+To see whether word order is related to the length of nouns and verbs we can first visualize the relationships by plotting the means of word class lengths (for arguments, predicates, nouns, verbs, etc) in relation to word order (SV, VS, free).
+Figure 3 shows the means of unique nouns and verbs (measure #1 above) in relation to word order, while Figure 4 shows the means of frequency-weighted nouns and verbs in relation to word order (measure #2 above).
+[^10] Additional statistical comparisons and plots are available at the linked repository (Fn 2).
+Observationally, the plots indicate that when frequency is taken into account, nouns are longer than verbs in SV languages but shorter than verbs in VS languages.
 [^11] It is also worth noting here that in this corpus words are overall shorter in SV languages than VS languages, which in turn are shorter than those in ‘free’ languages.
@@
 2020).
-Figure 3: Normalized lens of Args, Preds, by word order
-Figure 4: Frequency-weighted lens of Ns, Vs, by word order
-If we consider only nouns and only verbs, the findings are similar when we normalize the dataset by reducing the vocabulary to unique items.
-However, if we allow for frequency effects (i.e.
-we do not reduce the vocabulary to unique items) there is a clear correlation with word order: nouns are longer than verbs in SV languages, but shorter than verbs in VS languages (Figure 4; p < 0.001).
-This is a major finding that we will return to in the discussion.
-Setting it aside for now, we turn to two different tests of this measure relating to diachrony: predicting word order change, and the potential interaction with family membership (genealogy/descent) and area (geography).
+Figure 3: Normalized lengths of Nouns, Verbs, by word order
+Figure 4: Frequency-weighted lengths of Ns, Vs, by word order
+To verify this I conducted a repeated-measures ANOVA with the average lengths of nouns and verbs as the repeated measures and the word order (SV, VS, free) as the between subject factors.
+Results of the ANOVA indicate that the effect of frequency-weighted noun and verb lengths is highly significant (p < 0.001) for differentiating between word orders.
+However, the ANOVA does not account for potential effects of language relatedness or geographical area, a subject we turn to next.
+2.1.4 Controlling for genealogical and areal bias
+Given that human languages develop over time through interaction, it has long been recognized that the evolution of particular linguistic features can potentially be accounted for by descent from a common ancestral state or by contact between languages spoken in the same area (i.e.
+Galton’s problem, c.f.
+Bromham 2024).
+It is therefore common practice to account or control for such genealogical or areal effects in typological studies.
+While there are several proposals on the best way to handle this (ibid, see also extensive discussion in Guzmán Naranjo & Becker 2022), a basic method is to group observed languages according to their family membership and/or macroarea.
+In keeping with such approaches, for each language I identified family membership and linguistic area on the basis of Glottolog classifications, [^12] Glottolog (Hammarström et al.
+2024; https://www.glottolog.org) contains expert annotations for language family affiliations for over 7,000 languages, along with ‘macroarea’ identifiers and additional aggregated data from a variety of sources.
+For family membership I take the highest level node, based on the observation that all modern languages are the same age, and while their pathways/histories may differ, this does not impact whether they can be considered part of the same phylogenetic tree.
+filling in missing data via Ethnologue and other sources, and included them as grouping criteria in a linear mixed-effects regression model.
+From the dataset of 1,942 languages I examined four word-length-related continuous features to identify potential effects of genealogy and area on word order.
+The first two features are the frequency-weighted average length of nouns and verbs in a corpus.
+The third feature is the difference in length between nouns and verbs, which is related to minimization of complexity, based on the idea that a larger difference between lengths of word classes may effect the likelihood of that particular class being placed first.
+The fourth feature is the ratio of noun and verb lengths, and is related to maximization of complexity (via “entropy”).
+Here a lower ratio indicates that the particular word classes are less likely to be easily differentiable (and therefore predictable) based on their length.
+For each feature I conducted separate linear regressions, treating both grouping factors (family membership, macroarea) as random effects with word order (SV, VS, free) as the dependent variable and the feature as the independent variable (Table 1).
+[^13] Tables in this section use the Mixed Linear Model from the Python statsmodels library, and all features are scaled with the MinMaxScaler from scikit-learn (which scales values to between 0 and 1).
+Other approaches can easily be explored using code at the taggedPBC Github repository.
+Each feature was found to be significant, with little effect from either family membership or macroarea.
+Treating family and macroarea as nested within each other did not change this result.
+Number of Languages: 1,942
+iv Coef. Std.Err. z P>|z|
+NVlensdiff 0.281 0.094 2.996 0.003
+NVlensratio 0.255 0.080 3.174 0.002
+Nlen_freq 0.433 0.109 3.957 0.000
+Vlen_freq 0.499 0.100 4.998 0.000
+Table 1: Word length features in relation to word order (SV, VS, free) with family/area factors (for 1,942 languages)
+As an additional check (Table 2), I filtered the dataset based on number of languages in each family, first excluding single languages and isolates (leaving 1,864 languages), then excluding families with fewer than 10 members (leaving 1,653 languages).
+In each case all features remained significant, with little effect from family membership or macroarea.
+Number of Languages: 1,864
+iv Coef. Std.Err. z P>|z|
+NVlensdiff 0.321 0.095 3.360 0.001
+NVlensratio 0.283 0.081 3.478 0.001
+Nlen_freq 0.448 0.111 4.045 0.000
+Vlen_freq 0.539 0.101 5.318 0.000
+Number of Languages: 1,653
+iv Coef. Std.Err. z P>|z|
+NVlensdiff 0.382 0.101 3.766 0.000
+NVlensratio 0.323 0.086 3.751 0.000
+Nlen_freq 0.332 0.122 2.710 0.007
+Vlen_freq 0.475 0.108 4.394 0.000
+Table 2: Word length features in relation to word order (SV, VS, free) with family/area factors (for 1,864 and 1,653 languages respectively)
+It is worth noting here that coefficients of the independent variables are in most cases not above 0.5.
+However, when we see that these coefficients are combined with z-values above 2.0 and highly significant P-values (all below 0.01), it indicates a robust effect (Luke 2017) for each feature.
+It is also important to observe that two of the features are derived from the others, which means there is likely to be high collinearity, such that combining their values in a regression results in the most robust feature (’Vlen_freq’ in this case) being found to be significant while the others are not.
+That features are found to be collinear does not mean that they should be excluded from a model, however - features that point in the same direction can still add extra information, which can be useful in other tasks such as classification, to which we turn next.
+These observations from linear regressions are simply one way of looking at the data, highlighting the importance of word length in relation to word order.
+By including family and macroarea as independent random effects the results indicate that this relationship is not biased by either factor.
 2.2 Study 2: Predictive validity of word length
@@
 If we are strictly interested in identifying basic word order from corpora, it is clear that the N1 ratio already allows us to do so.
-However, we are also interested in understanding how noun and verb lengths may be related to word order.
-Being able to predict word order only from noun and verb lengths in a corpus for both a modern language and its ancestral language would indicate that such properties are related to how word order evolves over time.
-To investigate this question, for Study 2 I sourced additional corpora of tagged data from two pairs of historical and modern languages in which the modern language changed its word order:
+Study 1 has shown that word length features derived from corpora are also highly significant for differentiating word orders.
+This is relevant because we are interested in understanding how noun and verb lengths may impact how word order changes over time.
+Given two sets of corpora from two timepoints of a language, being able to predict word order from noun and verb lengths in a corpus for both the modern language and its ancestral language would indicate that such properties are related to how word order evolves.
+To investigate whether this is the case, for Study 2 I sourced additional corpora of tagged data from two pairs of historical and modern languages in which the the descendent changed its word order from the ancestral state:
   1. 1.
@@
 In the case of Classical Arabic I used the Quranic Arabic Corpus [^16] Version 0.4: https://corpus.quran.com and for Egyptian Arabic I used the BOLT Egyptian Arabic Treebank (Maamouri et al.
-2018).
-[^14] https://catalog.ldc.upenn.edu/LDC2018T23 (Discussion Forum) If the lengths of nouns and verbs in these corpora allows us to correctly classify Ancient Hebrew and Classical Arabic as VS and the modern varieties as SV, this would suggest that these are salient properties affecting how word order has developed diachronically.
-Importantly, given the textual differences between the taggedPBC (New Testament texts) and the test corpora (Biblical, Quranic, modern, and informal texts), such a finding would be significant.
+2018), [^17] https://catalog.ldc.upenn.edu/LDC2018T23 (Discussion Forum) which are also hand-tagged for parts of speech.
+If the lengths of nouns and verbs in these corpora allows us to correctly classify Ancient Hebrew and Classical Arabic as VS and the modern varieties as SV, this would suggest that these are salient properties affecting how word order has developed diachronically.
+Given the textual differences between the taggedPBC (New Testament texts) and the test corpora (Biblical, Quranic, modern, and informal texts), such a finding would be quite striking.
 In order to conduct this test, the data for Ancient Hebrew (ISO 639-3: hbo), Modern Hebrew (ISO 639-3: heb), Classical Arabic (cla) [^18] Classical Arabic does not have an assigned ISO 639-3 code or Glottocode.
@@
 and Egyptian Arabic (ISO 639-3: arz) were processed and the respective word classes were counted and measured as per the method outlined for Study 1 (each corpus > 400 sentences, see Table 3).
-The average length of nouns and verbs (including frequency effects) were then used as continuous features for a Gaussian Naive Bayes (GNB) classifier.
 Language ISO 639-3 Data Source Sentences Unique nouns Unique verbs
@@
 Table 3: Corpora used for testing historical word order change
-Three additional features were also used to expand the feature dimensions.
-The first is a categorical feature: whether a language’s nouns are longer than its verbs.
-The second is a continuous feature: the difference in length between nouns and verbs.
-This second measure is related to minimization of complexity, based on the idea that a larger difference between lengths of word classes may effect the likelihood of that particular class being placed first.
-The third measure is the ratio of noun and verb lengths, and is related to maximization of complexity (via “entropy”), whereby a lower ratio indicates that the particular word classes are less likely to be easily differentiable (and therefore predictable) based on their length.
-Using these five features (nounlen, verblen, nlonger, nlen-vlen, nlen/vlen), a GNB classifier was trained on the original combined word order dataset from the taggedPBC containing hand-coded classification of 1,150 languages (Modern Hebrew and Modern Arabic were removed), and then used to predict word order for our pairs of languages: Ancient Hebrew and Modern Hebrew on the one hand, Classical Arabic and Egyptian Arabic on the other.
+Five features related to word length were used to train a Gaussian Naive Bayes (GNB) classifier, namely the four features described in §2.1.4 above as well as a categorical feature: whether a language’s nouns are longer than its verbs.
+Using these five features (nounlen, verblen, nlen-vlen, nlen/vlen, nlonger), a GNB classifier was trained on the original combined word order dataset from the taggedPBC containing hand-coded classification of 1,150 languages (Modern Hebrew and Modern Arabic were removed).
+[^19] As noted previously, since we are concerned here with classification, the collinearity of relevant features is not an issue.
+The trained classifier was then used to predict word order for our pairs of languages: Ancient Hebrew and Modern Hebrew on the one hand, Classical Arabic and Egyptian Arabic on the other.
 Ancient Hebrew was classified as VS while Modern Hebrew was classified as SV by the trained classifier.
 Classical Arabic was classified as VS while Egyptian Arabic was classified as SV.
-As an additional check, I sourced data from another historical language, Old Irish, as represented in the Codex Paulinus Wirziburgensis, whereby Latin text of St.
+To double check this finding, I sourced data from another historical language, Old Irish, as represented in the Codex Paulinus Wirziburgensis, whereby Latin text of St.
 Paul’s epistles was annotated by Irish scholars during the middle of the eighth century.
 The dataset (Table 4; Doyle 2018) contains over 3,600 Irish glosses/sentences.
-As an Indo-European language, Old Irish is widely surmised to have changed its word order to verb-initial from an earlier state, so this check allows us to verify whether the same properties hold in relation to lengths of nouns and verbs.
+As an Indo-European language, Old Irish is widely surmised to have changed its word order to verb-initial from an earlier state (c.f.
+Watkins 1963; Hickey 2002), so this check allows us to verify whether the same properties hold in relation to lengths of nouns and verbs.
+I then removed Old Irish from the training dataset, retrained the classifier, and predicted on the features of the Würzburg Glosses.
 Indeed, the classifier accurately classified Old Irish as VS, based on the data extracted from the Würzburg Glosses.
@@
 Table 4: Irish corpus used for testing historical word order change
-Study 2 thus indicates that the relative lengths of nouns and verbs, when frequency is taken into account, can predict word order from corpora unrelated to the type of corpora contained in the taggedPBC, with various implications for our understanding of language processing and language change.
-This is a striking finding, and I am unaware of any other research that has been able to show predictive validity of a model related to word order change.
+Study 2 thus indicates that the frequency-weighted relative lengths of nouns and verbs can predict word order from corpora unrelated to the type of corpora contained in the taggedPBC, with various implications for our understanding of language processing and language change.
+This is a notable finding, and I am unaware of any other research that has been able to predict word order change based on measurement of corpus-based features.
 However, it is possible that this result is conditioned to some degree by the fact that the majority of the test cases are Semitic languages spoken in the same geographical area, so we now turn to an investigation of whether genealogical descent or location provides a better explanation for the correlation of noun and verb lengths with word order.
 2.3 Study 3: Testing variance predicted by word length
+The preceding studies have shown, first, that features related to word length differentiate between intransitive word orders crosslinguistically (Study 1) and, second, that such features can predict intransitive word order from corpora for both historical and modern languages (Study 2).
+While Study 1 controlled for the effect of language family and macroarea on word length features, Study 3 asks a slightly different question.
+That is: of the factors known to explain word order patterns in languages (family, area, word length), which is more explanatory?
 Dunn et al.
@@
 Although Jäger & Wahle (2021) counter this claim with additional data, we can also use the current dataset to see if descent is more explanatory of word order for a given language than the lengths of nouns and verbs.
-Additionally, in more recent work, Guzmán Naranjo & Becker (2022) show that areal affects are as important as family membership in predicting linguistic phenomena, finding that the results in Dunn et al can be just as easily predicted by language area.
-This is also a claim that we can test with our dataset.
-Accordingly, Study 3 can proceed with several (motivated) simplifying observations.
-First, for each language we can take the highest level node of family membership that is known based on current scholarship (via Glottolog).
-[^16] Glottolog (Hammarström et al.
-2024; https://www.glottolog.org) contains expert annotations for language family affiliations for over 7,000 languages, along with additional aggregated data from a variety of sources.
-This is because all modern languages are the same age, and while their pathways/histories may differ, this does not impact whether they can be considered part of the same phylogenetic tree.
-That is, if descent is more explanatory than word class length, we would expect languages belonging to the same tree to have a higher likelihood of sharing the same word order than not.
-Second, we can use the macroareas identified by Glottolog as well as the actual GPS coordinates of languages to control for areal affects.
-Third, we can run various analyses with differing sets of data in order to see how this affects the statistical robustness of the findings.
+Regarding language area, Guzmán Naranjo & Becker (2022) show that areal affects are as important as family membership in predicting linguistic phenomena, and that the results in Dunn et al.
+can be just as easily predicted by language area.
+This is also a claim that we can test with the taggedPBC dataset.
+To implement the comparison and identify the relative importance of known predictors we can use hierarchical linear regression (HLR).
+This approach allows researchers to observe the relative impact of particular features that are theoretically relevant to the question at hand by seeing how they influence the performance of successive models.
+In the case of word order, we can train a base regression model with a feature known to be relevant to word order (the N1 ratio), and then subsequent models that each incorporate additional features of interest (“language area/location”, “family membership”, “noun/verb lengths”).
+We can then compare the models to see whether the additional features increase the variance explained in the data via the F-value: an increase in F-value (change) and the P-value of this change indicates more/additional variance explained by the particular feature added in a model.
+Importantly, because we are not concerned with the significance of included features, each individual model can contain multiple (potentially collinear) features related to a given theoretical construct, as what is being measured is simply degree and significance of the change in fit of each model relative to the previous one.
+This approach also means we can run multiple analyses with differing sets of data in order to see how this affects the fit of the models.
 It may be, for example, that using a smaller sample gives a different result than a larger or more diverse one.
-Here, the more languages we have for a language family or geographical location, the higher the likelihood that any relationship identified is not due to chance.
-Finally, we can remove languages with “free” word order from consideration, as this leaves us with a binary distinction in word order (SV, VS) that opens up more possibilities for statistical tools.
-In particular, a binary dependent variable for word order means we can use hierarchical linear regression (HLR), which allows us to observe the relative impact of particular features that are theoretically relevant to the question at hand by seeing how they influence the performance of successive models.
-In the case of word order, we can train a regression model with a feature known to be relevant to word order (the N1 ratio), and then subsequent models that incorporate additional features of interest (“language area/location”, “family membership”, “noun/verb lengths”).
-We can then compare the models to see whether the additional features increase the variance explained in the data via the F-value: an increase in F-value (change) and the significance of this change indicates more/additional variance explained by the particular feature added in a model.
-Accordingly, for all languages in the taggedPBC I retrieved the family membership and geographical information from Glottolog, filling in missing data via Ethnologue and other sources.
-Then I filtered the data on four conditions: a) languages found in Dunn et al that are also present in the taggedPBC (120 languages), b) languages found in the taggedPBC that are in the four families investigated by Dunn et al (662 languages), c) languages in families with more than 75 members in the taggedPBC (1182 languages), and d) languages in families with 2 or more members in the taggedPBC (1,805 languages).
+Here, the more languages we have for a language family or linguistic area, the higher the likelihood that any relationship identified is not due to chance.
+One caveat of this approach is that when the dependent variable is categorical, an HLR requires that it be transformed to a single distinction.
+Thus in the ensuing observations I remove languages with “free” word order from consideration, as this leaves us with a binary distinction in word order (SV, VS) compatible with the HLR approach.
+Accordingly, for all languages in the taggedPBC I identified family membership and macroarea as noted in §2.1.4.
+Then, after removing the 63 languages with “free” word order, I filtered the data on four conditions: a) languages found in Dunn et al.
+that are also present in the taggedPBC (120 languages), b) languages found in the taggedPBC that are in the four families investigated by Dunn et al.
+(662 languages), c) languages in families with more than 75 members in the taggedPBC (1,182 languages), and d) languages in families with 2 or more members in the taggedPBC (1,805 languages).
 For each of these conditions, I ran an HLR, adding features of interest progressively to see how they impacted the fit of the successive models and whether particular features account for additional variance.
 The base model (Model 1) uses the N1 ratio as the primary feature for classification.
-This is because we know from comparison with expert determinations of word order (see Ring 2025) that the N1 ratio significantly correlates with the classification of languages as SV or VS.
-Due to the strong relationship observed between language area and typological properties of language (see Guzmán Naranjo & Becker 2022), Model 2 then includes GPS coordinates and macroarea identifiers for each language as features.
-Model 3 additionally incorporates descent (language family membership) as a feature (as per Dunn et al.
-2011), and Model 4 includes the average lengths of nouns and verbs (as per the relationship identified in this paper).
+This is because we know from comparison with expert determinations of word order (see Ring 2026) that the N1 ratio significantly correlates with the classification of languages as SV or VS.
+Due to the strong relationship observed between language area and typological properties of language (see Guzmán Naranjo & Becker 2022), Model 2 then includes macroarea identifiers for each language as features.
+Model 3 additionally incorporates descent (language family membership) as a feature (per Dunn et al.
+2011), and Model 4 includes the average lengths of nouns and verbs (per the relationship identified in this paper).
+[^20] In these models, continuous features (N1 ratio, N/V len) were scaled using the MinMaxScaler from scikit-learn, and categorical features (family, macroarea) were converted to numeric values.
 If descent is more explanatory than the lengths of nouns/verbs, this should be clear from the differences between the respective models for each condition.
@@
 a) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
-1 (N1 ratio) 120 45.27 6.4e-10
-2 (Lg area) 120 24.29 1.3e-14 12.78 2.87e-07 **
-3 (Family) 120 20.02 2.7e-14 2.05 0.154
-4 (N/V len) 120 15.75 2.9e-14 3.17 0.045 *
+1 (N1 ratio) 120 45.28 6.5e-10
+2 (Lg area) 120 46.63 1.28e-15 34.95 3.41e-08 **
+3 (Family) 120 30.86 9.65e-15 0.07 0.79
+4 (N/V len) 120 19.57 4.92e-14 1.91 0.15
 b) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
-1 (N1 ratio) 662 189.99 3.6e-38
-2 (Lg area) 662 95.65 4.1e-64 50.07 3.7e-29 **
-3 (Family) 662 78.23 2.4e-64 5.79 0.016 *
-4 (N/V len) 662 65.43 2.6e-71 21.30 1.0e-09 **
+1 (N1 ratio) 662 190.0 3.61e-38
+2 (Lg area) 662 182.17 1.05e-63 135.59 1.25e-28 **
+3 (Family) 662 122.02 7.73e-63 1.47 0.23
+4 (N/V len) 662 82.34 4.44e-67 15.02 4.18e-07 **
 c) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
-1 (N1 ratio) 1182 378.83 2.1e-73
-2 (Lg area) 1182 210.38 3.3e-136 116.99 2.6e-66 **
-3 (Family) 1182 169.09 1.3e-135 2.72 0.099
-4 (N/V len) 1182 127.47 2.3e-139 14.03 9.4e-07 **
+1 (N1 ratio) 1182 378.83 2.14e-73
+2 (Lg area) 1182 327.01 1.05e-113 208.55 1.2e-43 **
+3 (Family) 1182 218.01 1.79e-112 0.37 0.54
+4 (N/V len) 1182 132.79 1.16e-111 3.55 0.03 *
 d) Model N (obs) F-val P-val (F) F-val change P-val (F-val change)
-1 (N1 ratio) 1805 535.41 6.1e-104
-2 (Lg area) 1805 178.89 3.8e-129 46.52 5.8e-29 **
-3 (Family) 1805 143.49 2.3e-128 1.66 0.197
-4 (N/V len) 1805 108.16 2.2e-132 14.46 5.8e-07 **
+1 (N1 ratio) 1805 535.42 6.17e-104
+2 (Lg area) 1805 327.44 4.99e-122 92.34 2.35e-21 **
+3 (Family) 1805 218.53 6.92e-121 0.78 0.38
+4 (N/V len) 1805 136.84 3.62e-123 10.76 2.27e-05 **
 Table 5: Hierarchical Linear Regressions under 4 conditions
-From these analyses we can see that, as expected, the N1 ratio significantly predicts word order under all conditions.
+From these results we can see that, as expected, the N1 ratio significantly predicts word order under all conditions.
 Additionally, language area is a significant predictor of word order that adds to the strength of the classification.
 Language family (descent) does not add additional predictive power to our model under most conditions (over language area), whereas the length of nouns and verbs does.
-This analysis shows that the length of nouns and verbs in a language accounts for more variance and at the same time is more significant than that language’s family membership in determining its word order classification, under all test conditions.
-When we remove language area from the linear regression this is still the case, [^17] See additional statistical results at the linked repositories (Fn 2; Fn 5).
-indicating that while language area is a stronger predictor than family membership, lengths of nouns and verbs is yet a stronger one.
+This analysis indicates that the length of nouns and verbs in a language accounts for more variance and at the same time is more significant than that language’s family membership in determining its word order classification, under all test conditions.
+When we remove language area from the linear regression, [^21] See additional statistical results at the linked repositories (Fn 2; Fn 5).
+we find that family becomes a significant predictor of word order for a smaller sample of languages, but that under these conditions the length of nouns and verbs still accounts for additional variance and is also highly significant.
+Under conditions (c) and (d), with larger and more diverses samples, the model incorporating language family provides little explanation and is non-significant, while the subsequent model incorporating length of nouns and verbs explains additional variance that is highly significant.
 This largely aligns with Jäger and Wahl’s (2021) finding regarding the (lower) correlation between descent and linguistic structures in comparison to other explanations, with the caveat that Jäger and Wahl included language isolates, which is likely to greatly diminish the affect of descent.
 If a large number of languages in a sample are not related to others, clearly descent can have only a minimal impact on word order.
-By including only language families with two or more members in the taggedPBC (which excludes 111 languages), I show that descent does not explain much variance above and beyond language location/area, in contrast with noun and verb lengths.
-This also suggests that language family/descent is highly correlated with linguistic area, which conforms with expectations related to human history and migration.
+By including only language families with two or more members in the taggedPBC (which excludes 74 languages classified as having either SV/VS order), I show that descent does not explain much variance above and beyond language location/area, in contrast with noun and verb lengths.
+The test where language area was removed altogether from the HLR models also suggests that language family/descent is highly correlated with linguistic area, which conforms with expectations related to human history and migration.
 Additionally, the amount of variance explained seems to depend on the number of languages (and language families) in the sample, highlighting the need for more data for languages from under-represented language families.
 3 Discussion
+The preceding investigations have shown that the average length of nouns and verbs, weighted by their frequency in a corpus, correlates with intransitive word order crosslinguistically.
+This correlation is strong enough that such word length features can predict word order for both historical and modern languages, and is a stronger signal for word order than either language family or language area.
+These findings have various implications for our understanding of language processing and how languages change under an agentic evolutionary model.
+In particular, the effect of frequency and its relation to processing suggests some possibilities for reconceptualizing a theory of language behavior.
+The fact that a clear relationship between word order and word length is not observed unless frequency is taken into account indicates that the availability of corpus data is required to investigate questions related to crosslinguistic grammatical patterns and how they might develop.
+Here, the ability to group classes of words for observation allows for more fine-grained analyses regarding such patterns.
+For example, when we group word classes into arguments (prototypical agents/actors as represented by nouns, pronouns, proper nouns) and predicates (prototypical actions/states as represented by verbs, auxiliaries) there is a different realization of length than if we solely consider nouns and verbs.
+In the former condition all arguments are shorter than verbs crosslinguistically, while in the latter condition the differences in length emerge.
 That arguments overall are shorter in languages of the taggedPBC, across all word orders, suggests that there are certain (non-Noun) arguments affecting these observations.
-Particularly, the effect of PRON seems quite likely, whereby a tendency for pronouns to be shorter crosslinguistically, along with a preference for pronouns as arguments, may impact the observation that arguments are shorter overall than predicates.
+While this has yet to be investigated in detail, the effect of PRON seems quite likely, whereby a tendency for pronouns to be shorter crosslinguistically, along with a preference for pronouns as arguments, may impact the observation that arguments are shorter overall than predicates.
 This pattern aligns with the processing claim that “heavier” words (or word classes in this case) occur later in a sentence, partly validating Hypothesis 1 of study 1.
@@
 These two pressures can be (and, in fact, have been) considered as two different kinds of “efficiency”: production efficiency vs communicative efficiency.
-Under this view, production efficiency is related to length of words (as per Wasow 2002, 2022; Hawkins 2014) - speakers use shorter words whenever they can to efficiently produce utterances (see also Mahowald et al.
+Here production efficiency is related to length of words (as per Wasow 2002, 2022; Hawkins 2014) - speakers use shorter words whenever they can to efficiently produce utterances (see also Mahowald et al.
 2013).
@@
 However, the effect of word length is clearly a major part of the explanation for why we can observe (and predict) such a difference in word order between Ancient and Modern Hebrew as well as Classical and Egyptian Arabic via their respective corpora.
+The importance of frequency to any account of language evolution is that frequency reflects usage.
+The most frequent items in a corpus are those that are most used by speakers of a language.
+As such, under a statistical learning paradigm, they are most likely to influence generalizations made by speakers/hearers based on their input.
+This means that any discussion of linguistic forms over time ought to consider frequency as a factor in how such forms change.
+As an example to illustrate the potential impact of this theory, I offer the following thought experiment involving a hypothetical language.
+This hypothetical language is limited to nouns and verbs, which the speakers use in their interactions, in goal-directed behavior such as making tools.
+Nouns and verbs are the same length, and the informational content of each word is simply based on their communicative impact in the interactions.
+As speakers use the nouns and verbs to communicate about their tool-making, there is no pressure from noun/verb lengths on order of constituents, with order being determined by the need to identify objects/actions in immediate interaction.
+Over time, as speakers innovate and additional information begins to get encoded (such as deixis), the information will become associated with either nouns or verbs, possibly resulting in additional morphology, and potential affixation.
+The relevant noun/verb associated with such morphology then becomes “heavier” and provides more information for the interaction.
+This makes them more likely to be placed at the beginning of an utterance, since that makes it easier to predict what follows.
+The information-rich noun/verb has great utility and therefore has high frequency, becoming notable and relevant to statistical inference by language learners, who may then generalize an ordering rule that places the heavier constituent at the front of the speech stream.
+How does this relate to language change?
+In this thought experiment, we could consider what might happen if a language that had developed markers on verbs only (and therefore had VS order) innovated a new set of tools for particular tasks, but with only a single word for “tool”.
+As the tools became more prevalent, usage increased in interaction, and each tool began to acquire affixes that identified their specific use and other relations.
+This developed to the point that a listener could infer, from the named object, a particular context and subsequent utterances.
+At this point the nouns have become “heavier” than the verbs, with the resulting stronger pressure to place them before verbs in the speech stream, and the language would realize SV order.
+While this is only a thought experiment, it does align with what we know about the development of pidgins and creoles.
+Pidgins are languages that arise out of contact situations and tend to lack morphology.
+These can develop into creoles as they acquire native speakers (children who grow up speaking the pidgin), and they begin to exhibit increased morphological complexity as the younger speakers innovate forms.
+The possibility laid out in this thought experiment is also supported by the finding in this paper that the length of particular word classes can predict the basic word order of a language.
+In turn, this leads to some general predictions: if, on average, nouns are longer than verbs in a language, it is highly likely to have SV word order.
+If, on the other hand, verbs are longer, it is highly likely to have VS word order.
+The effect of frequency highlights the importance of conducting large-scale corpus based studies to investigate this and related concerns.
+Since languages are used by speakers and such speakers use certain words more than others, we can only observe this frequency effect (or “weight”) of word class length via corpora.
 The existence of multiple competing pressures affecting the development of language, as described by this Min-Max theory, is supported by work in clinical psychology (e.g.
@@
 Further, the data shows that all languages have a relatively high proportion of sentences in which the argument occurs before the predicate.
-In VS languages, the proportion of argument-initial to predicate-initial sentences is 1/2 or more (min: 0.56 n/v, 0.65 a/p), highlighting that even VS languages still place the noun/argument first in many constructions.
+In VS languages, the proportion of argument-initial to predicate-initial sentences is 1/2 or more (min: 0.56 n/v, 0.65 a/p), highlighting that even VS languages still place the noun/argument first in many constructions, and suggesting that there is a strong efficiency pressure toward SV order.
 Similarly, while all SV languages have a high proportion of N1 sentences (max 2.43 n/v, 7.17 a/p), the number of V1 constructions is non-zero, supporting a gradient approach to word order (Levshina et al.
 2023).
-The fact that the length of particular word classes can predict the basic word order of a language is a novel finding, leading to some general predictions: if, on average, nouns are longer than verbs in a language, it is highly likely to have SV word order.
-If, on the other hand, verbs are longer, it is highly likely to have VS word order.
-The effect of frequency cannot be ignored either, highlighting the importance of conducting large-scale corpus based studies to investigate this question.
-Since languages are used by speakers and such speakers use certain words more than others, we can only observe this frequency effect (or “weight”) of word class length via corpora.
 The idea that speakers balance multiple considerations in production is not new (Pike 1967; Jackendoff & Pinker 2005; Seržant & Moroz 2022; Hörberg & Sjons 2023) but what is novel is the idea that all humans experience a specific universal pressure to balance production/processing efficiency with information content.
@@
 Following on from Zipf’s and subsequent observations, there is no doubt that the length of strings is an important factor in processing streams, but clearly not all strings are created or treated equally by speakers.
-The studies presented here show that nouns are treated differently (or at least have a different distributional relationship) than other kinds of arguments, which leads to different realizations of word order crosslinguistically.
+The studies presented here show that nouns are treated differently (or at least have a different distributional relationship) than other kinds of arguments, which may then lead to different realizations of word order crosslinguistically.
 The pressures identified in this paper have wide-ranging implications and applications for any conceptualization of how language operates, whether in terms of understanding the intricacies of individual linguistic scenarios, designing intelligent systems that use language, or developing clinical interventions to assist speakers in various ways.
@@
 Given the time consuming nature of this task and the need for expert annotators, this is not yet possible, but ongoing research could address this by sourcing experts to annotate the data.
-Although I have shown (in Ring 2025 and via the linked repository) that the translation/alignment method does correspond well to hand-tagged data, it may be that this could be improved by tagging a small amount of data for each language in the PBC in order to train POS taggers that perform better than the current translation method.
+Although I have shown (in Ring 2026 and via the linked repository) that the translation/alignment method does correspond well to hand-tagged data, it may be that this could be improved by tagging a small amount of data for each language in the PBC in order to train POS taggers that perform better than the current translation method.
 Another limitation is that the dataset used, while exceedingly large in comparison to previous efforts, still only represents about a quarter of the world’s languages.